### v4

# Aperiodic points for outer billiards

Abstract.
Euclidean outer billiard on a regular polygon (that is not a triangle, square or a hexagon) has aperiodic points, i.e., points where all iterates of the outer billiard map are defined and yield pairwise distinct images. This result answers a question of R. Schwartz posed at ICM 2022.
The fourth named author has been partially supported by the HSE University Basic Research Program. Theorem 4.13 was supported by the Russian Science Foundation under grant no. 22-11-00177.
1. Introduction
Fix an orientation of the real affine plane $\mathbb{R}^{2}$. Then, the two half-planes bounded by a given oriented line can be uniquely marked as the left and the right half-planes, that is, as lying on the left and, respectively, on the right of the line. If $a$ and $b$ are two different points in the plane, then the line $ab$ will mean the line passing through $a$, $b$ and oriented from $a$ to $b$. Say that a point lies on the left (resp., on the right) of the line $ab$ if it lies in the left (resp., right) half-plane bounded by $ab$.
Figure 1. Outer billiard map on a regular pentagon $\Pi$. Here, the vertices of $\Pi$ are the roots of unity $\zeta^{k}$, where $\zeta=\exp(2\pi\mathbf{i}/5)$. The domain $\mathrm{dom}(f)$ is the union of the 5 sectors $V_{k}$, and the map $f_{\Pi}$ acts on $V_{k}$ as the half-turn about $\zeta^{k}$.
Consider an open convex polygon $\Pi$ in $\mathbb{R}^{2}$. The right (resp., left) outer billiard on $\Pi$ is the map $f_{\Pi}$ defined for points $z\in\mathbb{R}^{2}$ such that, for some vertex $a$ of $\Pi$, all other vertices of $\Pi$ lie on the left (resp., right) of $za$. For such points $z$, one sets $f_{\Pi}(z)=2a-z$, that is, $f_{\Pi}$ acts on $z$ as the central reflection (=a half-turn) about the vertex $a$. It is clear that the domain $\mathrm{dom}(f_{\Pi})$ of $f_{\Pi}$ is the complement in $\mathbb{R}^{2}$ of the closure $\overline{\Pi}$ of $\Pi$ united with finitely many rays, each emanating from a vertex of $\Pi$ and lying on a support line of $\Pi$. We are concerned with outer billiards on regular $N$-gons with integer $N>2$. In the sequel, $f_{\Pi}$ will stand for the right outer billiard on $\Pi$. See Fig. 1 for an illustration.
Recall that a point $z\in\mathbb{R}^{2}$ is called aperiodic for $f_{\Pi}$ if all iterates $f^{n}_{\Pi}(z)$ are defined for $z$ (where $n$ runs through all integers, positive or negative) and are all pairwise distinct. Our main result is the following.
Theorem 1.1.
Outer billiard on any regular $N$-gon with $N>4$ and $N\neq 6$ has aperiodic points.
Theorem 1.1 is proved by rather general methods, which may be applicable to wider classes of piecewise isometries.
1.1. Overview and related works
More generally, outer billiard can be defined for any compact convex figure in the plane (see [Tab93, DT05, Tab05] for overviews). First definition of outer billiard maps may be due to M. Day [Day47]; it reappeared in B. Neumann [Neu59]. J. Moser [Mos78] suggested outer billiard as a crude model of planetary motion: a simplification, in which the stability problem still has subtleties of the original context. Dynamical stability of outer billiard means that all orbits are bounded. This is the case when the figure is strictly convex and has sufficiently smooth boundary [Mos01, Dou82]. It is also the case [VS87, Koł89, GS92] for the so called quasi-rational polygons, which include all lattice polygons as well as all regular polygons. See also the thesis of D. Genin [Gen05] for the boundedness of orbits in a family of trapezoidal outer billiards. However, as was discovered in [Sch07], outer billiards on convex quadrilaterals can have unbounded orbits ([Sch09, Sch19] study these orbits in more detail).
Polygonal outer billiards are arguably the principal examples of Euclidean piecewise rotations (see [Goe00, Goe03] for overviews of piecewise isometries, including piecewise rotations). Piecewise rotations are a natural generalization of interval exchange transformations. They also found applications in electrical engineering, in particular, in the study of digital filters [Ash96, ACP97, CL88, CL90, Dav92, KWC96, Ogo92]. In [YFWC10], a sufficient condition for the existence of periodic points is given for a general class of piecewise isometries.
Previously known results concerning outer billiards on regular polygons are as follows. Apart from “trivial” (or “integrable”) cases of regular $N$-gons for $N=3,4,6$ (these “trivial” cases are characterised by the property that the group of isometries generated by the half-turns about all the vertices acts properly discontinuously on the plane), all understood cases possess dynamical self-similarities, i.e., Euclidean similarities that conjugate the first return map of some polygon with the first return map of a strictly smaller polygon. S. Tabachnikov [Tab95] described a self-similarity for $N=5$ and its implications for orbit coding and the Hausdorff dimension of the Julia set, which eventually led to a precise symbolic model in this case (Bedaride-Cassaigne [BC11]). G.H. Hughes [Hug12, Hug13, Hug16, Hug21] has a number of computer assisted experimental conjectures concerning the values of $N$ up to about 20. Cases $N=8,10,12$ are rigorously covered by the thesis of Ph. Rukhovich [Ruk22], which was based on earlier observations and conjectures concerning self-similarities of these systems as well as automated computations using precise number field arithmetics. Previous results of R. Schwartz [Sch14] embedded the case $N=8$ into a continuous family with an action of a suitably defined renormalization operator. Yet earlier, in [AKT01], a similar piecewise rotation of the torus was studied. Cases $N=7$, $9$ display multi-fractal structure (independent self-similarities leading to uncountable families of pairwise disjoint closed invariant subsets of aperiodic points), which may be suggested by experimental data of [Hug13]; a computer aided proof of these phenomena is a work in progress by V. Zgursky and the authors of this paper.
In his ICM 2022 address [Sch21], R. Schwartz stated Conjecture 6.5, which reads: “outer billiard on the regular $n$-gon has an aperiodic orbit if $n\neq 3,4,6$”. This paper settles the conjecture. Previously, E. Gutkin and N. Simanyi [GS92] posed a more general question: “Let $P$ be a quasi-rational but not rational polygon. We expect that the outer billiard about $P$ to have nonperiodic orbits.” Methods of this paper may also shed some light on this question but this is a plan for the future.
1.2. Methods
Even though the outer billiard map itself is an example of a piecewise rotation, the domain of this piecewise rotation is unbounded, which makes things more complicated. A remedy to this complication has been known for a long time: there are bounded regions invariant under $f_{\Pi}$; the restrictions of $f_{\Pi}$ to these regions are bounded piecewise rotations, some of which are easy to describe. Dynamics of $f_{\Pi}$ on a bounded invariant polygon can be easily reduced to that of a polygon exchange transformation, i.e., an invertible piecewise isometry acting by parallel translations on all components of its domain. Next, the main issue when looking for aperiodic points of a polygon exchange transformation of a bounded polygon is to rule out full periodicity of the map on the complement of finitely many lines.
For this problem — and here comes the principal novelty of this paper — we employ scissors congruence invariants. Two polygons are scissors congruent if one can cut one polygon into finitely many polygonal pieces and assemble the other polygon from these pieces. If $\mathsf{G}$ is a subgroup of $\mathrm{Isom}(\mathbb{R}^{2})$, the full isometry group of $\mathbb{R}^{2}$, whose elements are the only ones allowed to act on the pieces, then we speak of $\mathsf{G}$-scissors congruence. A case of special importance for us is $\mathsf{G}=\mathsf{T}$: the group of all translations. By ($\mathsf{G}$-)scissors congruence invariants, we mean functions on polygons that take the same values on $\mathsf{G}$-scissors congruent polygons. Dynamical versions of $\mathsf{T}$-scissors congruence invariants are conjugacy invariants of polygon exchange transformations. [^1] Polygon exchange transformations are themselves scissors congruences, and dynamic invariants are, in a sense, scissors congruence invariants of these scissors congruences.
Such invariants, in principle, can lead to very general necessary conditions of periodicity, in the form of vanishing of these invariants. A sample condition of this type follows from Theorem 2.29: for a periodic polygon exchange transformation, any dynamic invariant (as described in Section 2.6) must vanish. The principle idea of Theorem 1.1 is simple: compute suitable dynamic invariants and show they are non-zero. Once it is established that a bounded restriction of $f_{\Pi}$ is not periodic, establishing the existence of periodic points is not immediate, but this part is less conceptual and more technical.
Using earlier results (cf. [Ruk22]), the Main Theorem can be reduced to the case of even $N$: for an odd $N$, there is a certain correspondence between outer billiards on the $N$-gon and on the $2N$-gon, allowing to conclude that the existence of aperiodic points in one of these systems implies the same for the other. Thus, it suffices to assume that $N$ is even. We also need to reduce the dynamics of $f_{\Pi}$ to that of a polygon exchange transformation, i.e., a piecewise isometry acting by translations. One obvious choice is the second iterate $f^{2}_{\Pi}$, however, a different choice allows to significantly simplify the computations. Namely, consider the half-turn $M$ about the center of $\Pi$, and form the composition $g:=M\circ f_{\Pi}$. It acts as a piecewise translation, see Fig. 3. Also, an $f_{\Pi}$-invariant bounded domain is invariant under $M$ and hence under $g$, and the dynamics of $g$ is not much different from that of $f_{\Pi}$, since $f_{\Pi}$ commutes with $M$. For these reasons, if $g$ is not periodic, then neither is a bounded restriction of $f_{\Pi}$. To establish non-periodicity of $g$, we use a two-dimensional dynamic invariant based on the Hadwiger–Glur valuation [HG51]. Explicit computation of this invariant and verification that the result is nonzero are not hard.
Yet another approach to proving non-periodicity of a polygon exchange transformation is implemented in a special case of Theorem 1.1, under the additional assumption that $N$ is twice an odd number. However, this second approach is potentially applicable to wider classes of piecewise isometries, not necessarily piecewise translations. It reduces a two-dymensional piecewise isometry to a one-dimensional one, that is, to an interval exchange transformation called a boundary quasi-commutator. Establishing that the Sah–Arnoux–Fathi invariant of the latter is nonzero is enough for proving non-periodicity of the original 2D map.
Even though the two approaches are, by the end of the day, rely on similar ideas (non-vanishing of dynamic invariants), their implementations look different, and the invariants employed are also different. Any one of the two approaches is enough to prove the main theorem, however, we describe both. This is done on purpose: one or the other approch may work better for other problems.
1.3. Organization of the paper
Section 2 discusses general concepts associated with scissors congruences, including their dynamical aspects. As a special case, we consider polygon exchange transformations, which act as piecewise translations. Dynamic invariants of polytope exchange transformations of a given polygon are certain homomorphisms of the group of such transformations to the additive group of a real vector space. It follows that these dynamic invariants are conjugacy invariants under orientation preserving piecewise Euclidean conjugacies, and, for this reason, they are also called dynamic invariants. Dynamic Hadwiger invariants can be obtained as special cases of more general dynamic invariants; they are based on (translation) scissors congruence invariants of Hadwiger and Glur [HG51] initially discovered as a means of solving translation scissors congruence problem for planar polygons.
Assuming that $N$ is even, Section 3 deals with the outer billiard map $f_{\Pi}$ on a regular $N$-gon $\Pi$. There is a bounded domain $X$ invariant under $f_{\Pi}$; see Section 3.1. As explained above, the map $g_{X}:=M\circ f_{\Pi}|_{X}$ is a polygon exchange transformation of a bounded domain. Here $M$ is the half-turn about the center of $\Pi$. Computing the dynamic Hadwiger invariant of $g_{X}$ (Theorem 3.6) and establishing that it is nonzero (Corollary 3.7) yields that $f_{\Pi}$ is not periodic on $X$ (Theorem 3.8). We still need to upgrade this result to existence of aperiodic points, which is done in Section 6.
Further on, in Section 4, we explore a different approach. Boundary quasi-commutators are defined in Section 4.1 for arbirtary piecewise rotations of bounded polygons. In fact, with the same piecewise rotation $f$, one can associate many boundary quasi-commutators. However, if $f$ is periodic, then all of the corresponding quasi-commutators have vanishing SAF invariants, see Theorem 4.3. Next, in Section 4.2, we define induced billiard map $f_{\mathrm{ind}}$ associated with $f_{\Pi}$: basically, this is a quotient dynamical system, in which the rotations of $\Pi$ are quotiented out. When $N$ is twice odd, there is an invariant triangle $U$ for $f_{\mathrm{ind}}$, which makes computations simpler. For the restriction of $f_{\mathrm{ind}}$ to $U$, we compute the SAF invariant of the associated boundary quasi-commutator in Theorem 4.13. Our computation uses that a certain map of the boundary $\partial U$ of $U$ naturally associated with $f_{\mathrm{ind}}$ turns out to be an irrational rotation. Results of Section 4 give an alternative proof of Theorem 3.8 and are otherwise irrelevant for the proof of Theorem 1.1, even though these results are potentially useful for studying wider classes of Euclidean scissors congruences, not necessarily piecewise translations.
Section 5 is a digression that discusses general notions and results of topological dynamics, some of which are classical. These results are needed in order to extract a canonical invariant subset of the Julia set that would be perfect (hence uncountable) and contained in a finite graph consisting of boundary points.
The proof of Theorem 1.1 is completed in Section 6. Namely, the main theorem of that section, Theorem 6.1, states that nonvanishing dynamic Hadwiger invariants together with some minor additional assumptions on a polygon exchange transformation (automatically fullfilled in the case of outer billiards on regular polygons) imply the existence of aperiodic points. Even though this part of the proof is technically involved, it appears conceptually less important than Theorem 3.8.
Concluding Section 7 summarizes the arguments and discusses their potential role in further research.
1.4. Acknowledgements
Valery Zgursky was a member of the team at an earlier stage of the project and co-authored a previous version of the paper; with gratitude, we acknowledge his contribution to the overall understanding of the problem and intuition through a series of computer experiments. We are grateful to Victor Kleptsyn for fruitful discussions and for communicating a proof of Proposition 2.33, to Alexandra Skripchenko for consultations on the topic of interval exchange transformations.
2. Dynamics of scissors congruences
In this section, we introduce terminology related to scissors congruences and, specifically, polygon exchange transformations, which are two dimensional generalizations of interval exchange maps.
2.1. Basic terminology
Given a compact polygon $P$ in the plane, not necessarily convex, even not necessarily connected, one can cut it into finitely many polygonal pieces and form a new polygon $Q$ from these pieces. In this case, we talk about a scissors congruence between $P$ and $Q$ and call the polygons $P$ and $Q$ scissors congruent. A more precise definition specifies a precise meaning of “cutting into pieces” as well as describes what kind of motions are allowed on the pieces.
Definition 2.1 (Scissors congruence).
Fix a subgroup $\mathsf{G}$ in the full group of Euclidean isometries of the plane. Also, consider two open bounded polygons $P$ and $Q$ in the plane, not necessarily convex and not necessarily connected. Suppose that
$\overline{P}=\overline{P_{1}\cup\dots\cup P_{n}},\quad\overline{Q}=\overline{Q_{1}\cup\dots\cup Q_{n}},$
where $P_{i}$ and $Q_{i}$ are open convex polygons such that $g_{i}P_{i}=Q_{i}$ for some $g\in\mathsf{G}$, moreover, $P_{i}\cap P_{j}=Q_{i}\cap Q_{j}=\varnothing$ for $i\neq j$. In this case, $P$ and $Q$ are said to be $\mathsf{G}$-scissors congruent. The map $f$ defined on $P_{1}\cup\dots\cup P_{n}$ and acting as $g_{i}$ on each $P_{i}$ is called the corresponding ($\mathsf{G}$-)scissors congruence.
A scissors congruence $f$ between $P$ and $Q$ is therefore a partially defined map; write $\mathrm{dom}(f)$ for the domain of $f$, that is, for the union of $P_{i}$s. Clearly, the inverse map $f^{-1}$ is defined on $Q_{1}\cup\dots\cup Q_{n}$; it is a scissors congruence between $Q$ and $P$. The domain of $f^{-1}$ coincides with the image of $f$, that is $\mathrm{dom}(f^{-1})=f(\mathrm{dom}(f))$. Now assume that $P=Q$; in this case, $f$ generates a dynamical system. Given a positive integer $n$, write $f^{n}$ for the $n$-th iterate of $f$ and $f^{-n}$ for the $n$-th iterate of $f^{-1}$. Setting also $f^{0}=id$ makes iterates $f^{n}$ well defined scissors congruences for all integer values of $n$. Note that the domains of $f^{n}$ are, in general, different for different values of $n$.
Remark 2.2 (Groupoid structure).
Recall that a groupoid is by definition a small category, all of whose morphisms are isomorphisms. There is a natural groupoid, the scissors congruence groupoid of $\mathsf{G}$, whose objects are bounded polygons and whose arrows are classes of $\mathsf{G}$-scissors congruences under the following equivalence relation: $f_{1}\sim f_{2}$ if $f_{1}$ and $f_{2}$ become equal after a polygonal subdivision of their domains. In any groupoid, all automorphisms of an object form a group. Thus, it makes sense to talk about the group of all scissors automorphisms of a given bounded polygon.
Scissors automorphisms belong to a more general class of piecewise isometries discussed, e.g., in [Goe00]. We adopt standard terminology and notation related to partially defined maps. In particular, for a partially defined map $f$ of a set $P$ and a subset $A\subset P$, the image $f(A)$ is defined as $f(A\cap\mathrm{dom}(f))$. Given a scissors congruence of a polygon $P$ and a subpolygon $A\subset P$, say that $A$ is invariant under $f$ if $f(A)\subset\overline{A}$; then automatically $f^{-1}(A)\subset\overline{A}$ since the area is preserved.
Remark 2.3 (Connection with outer billiards).
Consider the outer billiard map $f_{\Pi}$ associated with a convex polygon $\Pi\subset\mathbb{R}^{2}$. We claim that $f_{\Pi}$ is a piecewise rotation of $\mathbb{R}^{2}\setminus\overline{\Pi}$, that is, $f$ acts on each component of $\mathbb{R}^{2}\setminus\overline{\Pi}$ as an orientation preserving Euclidean isometry. Indeed, label the vertices of $\Pi$ as $\zeta^{0}$, $\dots$, $\zeta^{N-1}$, and let $V_{i}$ be the locus of all points $z\in\mathbb{R}^{2}\setminus\overline{\Pi}$ such that $f_{\Pi}$ acts on $z$ as the reflection in $\zeta^{i}$. Then $V_{i}$ are convex cones (=sectors) bounded by continuations of the edges of $\Pi$. On each $V_{i}$, the outer billiard map acts as the rotation by $\pi$, that is, as a half-turn, with center $\zeta^{i}$. Note that the domain of $f$ is the union of all $V_{i}$, that is, the complement in $\mathbb{R}^{2}\setminus\overline{\Pi}$ and $N$ straight line rays emanating from the vertices of $\Pi$. A way of obtaining a scissors congruence out of $f_{\Pi}$ is restricting the map $f_{\Pi}$ to a suitably chosen invariant bounded polygon; this is indeed possible, as we will see later.
Next, we define a number of notions associated with scissors automorphisms. Since the whole area of research is rather new, and terminology still differs from author to author, we allow ourselves to introduce terminology which we find the most suitable.
Definition 2.4 (Pieces).
Let $f$ be a scissors automorphism of a polygon $\Delta$. (Forward) pieces of $f$ of level $n\geqslant 0$ are by definition components of $\mathrm{dom}(f^{n})$. By definition, the only level 0 piece of $f$ is $\Delta$. Similarly, backward pieces are defined as components of $\mathrm{dom}(f^{-n})$ for $n>0$. Finally, level $n>0$ bi-pieces are by definition components of $\mathrm{dom}(f^{n})\cap\mathrm{dom}(f^{-n})$. As it is easy to see, for every level $n$, there are finitely many level $n$ forward or backward pieces, or bi-pieces. If the level of pieces is not indicated, we always mean level one pieces.
As follows from this definition, if $P$ is any forward piece of $f$, and $Q$ is any component of $\mathrm{dom}(f^{-1})\cap P$, then $f^{-1}(Q)$ is also a forward piece of $f$. Similar properties hold for backward pieces.
Definition 2.5 (Boundary, periodic, aperiodic points).
Given a scissors isomorphism $f$ of $\Delta$, define a boundary point of $f$ as a point $z\in\overline{\Delta}$, for which there is $n\in\mathbb{Z}$ such that $z\notin\mathrm{dom}(f^{n})$, that is, some positive or negative iterate of $f$ is not defined at $z$ (which includes the case $z\in\partial\Delta$). Clearly, boundary points of $f$ are precisely boundary points of all forward or backward pieces of $f$. If $f^{n}(z)=z$ for some $n\neq 0$ (which means, in particular, that $z\in\mathrm{dom}(f^{n})$), then the point $z$ is called periodic under $f$. The smallest positive integer $n$ with this property is called the period of $z$. It is immediate from the definition that periodic points are not boundary points. Aperiodic points are defined as points of $\Delta$ that are neither periodic points nor boundary points.
See [Goe03] for an overview of some results and open problems in dynamics of piecewise isometries and, in particular, in dynamics of scissors automorphisms.
2.2. Fatou and Julia sets
The following notions are defined by analogy with holomorphic dynamical systems.
Definition 2.6 (Fatou and Julia sets).
Let $f$ be a scissors automorphism of a polygon $\Delta$. An open subset of the plane is said to be essential (for $f$) if it contains either $f$-periodic points of arbitrarily large period, or at least one $f$-aperiodic point. Define the Julia set $J=J(f)$ of $f$ as the set of all points $z\in\overline{\Delta}$ such that every open neighborhood of $z$ is essential. The complement of $J$ is called the Fatou set $F(f)$ of $f$. It coincides with the set of points that admit non-essential neighborhoods.
The Fatou set is cut by the set of boundary points into open domains which are either periodic or wandering (an open domain $\Omega\subset F(f)$ is said to be wandering if $f^{n}(\Omega)\cap\Omega=\varnothing$ unless $n=0$). The latter case is of course impossible if $\Delta$ has finite area.
Definition 2.7 (Periodicity).
A scissors automorphism $f$ of $\Delta$ is called periodic if $f^{n}=id_{\mathrm{dom}(f^{n})}$ for some integer $n>0$. The smallest such $n$ is called the period of $f$. Equivalently, $f$ is periodic of period $n$ if it represents an order $n$ element in the group of scissors automorphisms of $\Delta$.
From the dynamical viewpoint, periodic scissors automorphisms are not very interesting. On the other hand, given a scissors automorphism, it may be difficult to see whether it is periodic or not. This paper introduces some general tools for establishing non-periodicity.
It is clear from the definition that the Julia set $J$ is compact, but it may be empty. Elements of $J$ are also called Julia points.
Lemma 2.8.
Let $f$ be a scissors automorphism. The Julia set of $f$ is nonempty if and only if $f$ is non-periodic.
Proof.
Clearly, if $f$ is periodic, then there are no essential open sets, hence $J$ is empty. On the other hand, if $f$ is non-periodic, then there is either an aperiodic point, which necessarily lies in $J$, or a sequence of periodic points whose periods tend to infinity, and any partial limit of this sequence, i.e., the limit of any converging subsequence, is in $J$. ∎
If a Julia point is not a boundary point, then it is necessarily aperiodic. However, some Julia points may be boundary points. Recall that an aperiodic point $z$ is recurrent if, for any neighborhood $O$ of $z$, there is a positive interger $n>0$ with $f^{n}(z)\in O$.
Lemma 2.9.
Any essential open set contains either periodic points of arbitrarily large period or a recurrent aperiodic point.
Proof.
Recall that $f$ preserves the Euclidean area, in particular, the Poincaré recurrence theorem is applicable to $f$. It says that all points of $\Delta$ are recurrent, except possibly for a set of zero area.
Let $O$ be an essential open set. Suppose first that it does not contain periodic points. Then almost all points of $O$ are aperiodic, and, by the Poincaré recurrence theorem, almost all of them are recurrent. We may now assume that every open subset of $O$ contains periodic points — otherwise, it has recurrent aperiodic points, and we are done. For any integer $n>0$, at least one component $O_{n}$ of $O\cap\mathrm{dom}(f^{n})$ is essential. Since $O_{n}$ is essential, it is not periodic, in particular, there are periodic points in $O_{n}$ of period $>n$. ∎
Given a positive integer $n$, set $\Gamma_{n}:=\overline{\Delta}\setminus\mathrm{dom}(f^{n})$ and call $\Gamma_{n}$ the boundary graph of level $n$. One also defines boundary graphs of negative levels by setting $\Gamma_{-n}:=\overline{\Delta}\setminus\mathrm{dom}(f^{-n})$. These are indeed plane graphs whose edges are straight line segments. Our convension is that $f^{0}=id$ and $\mathrm{dom}(f^{0})=\Delta$, hence we set $\Gamma_{0}$ to be the boundary $\partial\Delta$ of $\Delta$. Note that $\Gamma_{n}\subset\Gamma_{m}$ if both $n$, $m$ are positive or both are negative, and $|n|<|m|$. Complementary components of $\Gamma_{n}$ for $n>0$ are precisely level $n$ forward pieces.
Define the level function $\mathrm{lev}:\overline{\Delta}\to\mathbb{Z}_{\geqslant 0}\cup\{\infty\}$ as the function that takes value 00 on $\partial\Delta$, value $n$ on $\Gamma_{n}\setminus\Gamma_{n-1}$ for every $n>0$, and value $\infty$ at all other points of $\Delta$. It follows from the definition that the set $\Gamma_{n}$ can be given by the inequality $\mathrm{lev}\leqslant n$. Also, the latter property determines the function $\mathrm{lev}$ uniquely. The function $\mathrm{lev}_{-1}$ is by definition the level function for $f^{-1}$. Set $\mathrm{bilev}:=\min(\mathrm{lev},\mathrm{lev}_{-1})$ so that the set defined by the inequality $\mathrm{bilev}\leqslant n$ is precisely $\Gamma_{n}\cup\Gamma_{-n}$. Call $\mathrm{bilev}$ the bi-level function.
Lemma 2.10.
The level function is lower semi-continuous: for every non-negative integer $n$, the set $\{z\in\overline{\Delta}\mid\mathrm{lev}(z)>n\}$ is open. Similar claim holds for the bi-level function.
Proof.
Indeed, the set $\{z\in\overline{\Delta}\mid\mathrm{lev}(z)>n\}$ coincides with $\mathrm{dom}(f^{n})$, that is, with the complement of $\Gamma_{n}$, which is a union of finitely many open convex polygons. The claim about the bi-level function is similar and is left to the reader. ∎
The full forward boundary $\Gamma_{+}=\Gamma_{+}(f)$ of $f$ is defined as the set of all points $z\in\overline{\Delta}$ with $\mathrm{lev}(z)<\infty$. In other words, $\Gamma_{+}$ is the union of all $\Gamma_{n}$ for $n\geqslant 0$. Similarly, the full backward boundary $\Gamma_{-}$ of $f$ can be defined as the union of $\Gamma_{-n}$ for all $n\geqslant 0$. Boundary points are precisely points of $\Gamma:=\Gamma_{+}\cup\Gamma_{-}$; the latter set is called the full boundary of $f$.
Lemma 2.11.
The map $f$ is periodic if and only if the level function takes only finitely many values, i.e., if and only if $\Gamma_{+}=\Gamma_{n_{*}}$ for some $n_{*}>0$. In the latter case, $\Gamma_{-}=\Gamma_{+}=\Gamma$.
Proof.
Suppose that $f$ is periodic of period $n_{*}$, then $\Gamma_{+}=\Gamma_{n_{*}}$. Conversely, assume that $\Gamma_{+}=\Gamma_{n_{*}}$ for some $n_{*}>0$. The complement of $\Gamma_{+}$ is forward invariant under $f$, hence all complementary components are periodic. It follows that the complement of $\Gamma_{+}$ is also backward invariant under $f$, that is, $\Gamma_{-}=\Gamma_{+}=\Gamma$, as claimed. ∎
2.3. Polygon exchange transformations
Let $\mathsf{T}$ be the group of all translations of the plane; $\mathsf{T}$-scissors automorphisms are also known as translation scissors congruences or as polygon exchange transformations. The notion of a polygon exchange transformation admits a straightforward generalization to higher dimensions, where one considers polytope exchange transformations. Dimension one polytope exchange transformations are called interval exchange transformations. An interval exchange transformation can be defined on a straight line interval or a finite union of intervals.
Given a polytope, i.e., a bounded polyhedron, $\Delta$ of dimension $d$, we write $\mathrm{PET}(\Delta)$ for the group of polytope exchange transformations of $\Delta$, that is, the group of all automorphisms of $\Delta$ in the $\mathsf{T}$-scissors congruence groupoid. More precisely, elements of $\mathrm{PET}(\Delta)$ are classes of polytope exchange transformations with respect to the equivalence relation identifying two maps if they coincide except on a finite union of affine hyperplanes. Periodic polytope exchange transformations yield, in particular, periodic elements of the group $\mathrm{PET}(\Delta)$. If $d=1$, then one writes $\mathrm{IET}(\Delta)$ for the group of all interval exchange transformations of $\Delta$. Abbreviations IET and PET stand, of course, for “interval exchange transformations” and “polytope exchange transformations”, respectively. We concentrate on dimensions two or one in this paper even though some notions and results make sense for arbitrary dimensions.
Example 2.12 (Interval swap maps).
Let $I$ be an interval and $I_{1}$, $I_{2}\subset I$ be disjoint (except possibly a shared endpoint) subintervals of equal length. Then there is an interval exchange transformation $s$ that takes $I_{1}$ onto $I_{2}$, takes $I_{2}$ onto $I_{1}$, and leaves all other points of $I$ fixed; of course, $s$ may remain undefined on the common endpoint of $I_{1}$ and $I_{2}$. Call $s$ an interval swap map.
Two important observations apply to polygon exchange transformations but not to more general piecewise rotations.
Lemma 2.13.
Any polygon exchange transformation has the following propery: for some positive integer $m_{*}$, at most $m_{*}$ pairwise disjoint pieces, no matter which levels, can share a boundary point.
Here, by pieces we mean forward or backward pieces, or bi-pieces.
Proof.
This is a simple geometric observation. Note that the graph $\Gamma_{1}$ lies in a union of, say, $m$ lines. Then there are at most $m$ lines such that the edges of all pieces are parallel to these lines. It follows that at most $2m$ pairwise disjoint pieces can share the same boundary point $z$: through $z$, draw all $\leqslant m$ lines parallel to the edges of $\Gamma_{1}$; these lines divide the plane into $\leqslant 2m$ sectors; disjoint pieces must lie in different sectors. ∎
A sequence of open polygons $Q_{n}$ indexed by nonnegative integers $n$ is said to be nested if $Q_{n+1}\subset Q_{n}$ for all $n\geqslant 0$. When writing a sequence, we use parentheses; thus $(Q_{n})$, or $(Q_{n})_{n}$ if one needs to indicate the index $n$, is a sequence indexed by $n$, and $Q_{n}$ are individual terms of this sequence.
Definition 2.14 (Nests and their impressions).
Say that a sequence $(Q_{n})$ is a nest for $f$, or an $f$-nest, if it is nested, and every $Q_{n}$ is a level $n$ bi-piece. Often, $f$ will be omitted from the notation of a nest, if it is clear from the context. The intersection of $\overline{Q}_{n}$ is called the impression of the nest $(Q_{n})$.
Another key observation, closely related to Lemma 2.13 and based on the same ideas, is the following:
Lemma 2.15.
A point of $\Delta$ can lie in the impressions of only finitely many nests, in fact, no more than $m_{*}$ nests, where $m_{*}$ is the same as in Lemma 2.13.
Proof.
By way of contradiction, assume that there are $m>m_{*}$ pairwise distinct nests whose impressions contain $z$. Any two distincts nests $(P_{n})$ and $(Q_{n})$ satisfy $P_{n}\cap Q_{n}=\varnothing$ for sufficiently large $n$. It follows that there is a large positive integer $n$ with the property that the level $n$ components in all the $m$ nests are pairwise disjoint. This contradicts Lemma 2.13. ∎
Lemma 2.13 also motivates the notion of an elementary sector.
Definition 2.16 (Elementary sectors).
An elementary sector at a point $z\in\overline{\Delta}$ is by definition a sector $\Lambda_{z}$ (i.e., a convex cone) with vertex $z$ and with the following two properties:
  1. (1)
    the germ of $\Lambda_{z}$ at $z$ coincides with a germ of some bi-piece;
  2. (2)
    any sector with property (1) includes $\Lambda_{z}$ or is disjoint from $\Lambda_{z}$.
Thus, elementary sectors are minimal possible sectors formed by bi-pieces.
Yet a third formulation, essentially repeating the content of Lemmas 2.13 and 2.15, is as follows.
Lemma 2.17.
Given any point in $\overline{\Delta}$, there are at most $m_{*}$ sectors at this point.
Here, $m_{*}$ is again the same as in Lemma 2.13. For example, if $z$ is not a boundary point of $f$, then it lies in the impression of only one nest (namely, the level $n$ component of this nest is recovered as the level $n$ bi-piece containing $z$), and the only elementary sector at $z$ is the entire plane.
2.4. Sah–Arnoux–Fathi invariants
We proceed by reviewing an invariant of an interval exchange transformation called the Sah–Arnoux–Fathi invariant (abbreviated as SAF invariant) [Arn81, Sah81]. Consider an interval exchange transformation $g$ defined on an interval $I$. Then the SAF invariant $\mathrm{Inv}(g)$ is an element of $\mathbb{R}\otimes_{\mathbb{Q}}\mathbb{R}$ defined as follows. If subintervals $I_{1}$, $\dots$, $I_{n}$ of lengths $\lambda_{1}$, $\dots$, $\lambda_{n}$, respectively, form a partition of $I$, and the restriction of $g$ to each $I_{i}$ is a translation by $\tau_{i}\in\mathbb{R}$, then
$\mathrm{Inv}(g)=\lambda_{1}\otimes\tau_{1}+\lambda_{2}\otimes\tau_{2}+\dots+\lambda_{n}\otimes\tau_{n}.$
The following theorem summarizes some (classical) important properties of the SAF invariant (a nice and comprehensive exposition of these results is contained in [Vor17]):
Theorem 2.18.
Every periodic interval exchange transformation has zero SAF invariant. Also, every interval swap map has zero SAF invariant. The map $g\mapsto\mathrm{Inv}(g)$ is a homomorphism from the group of all interval exchange transformations of a given interval to the additive group of $\mathbb{R}\otimes_{\mathbb{Q}}\mathbb{R}$; the image of this homomorphism lies in $\mathbb{R}\wedge_{\mathbb{Q}}\mathbb{R}$.
The map $\mathrm{Inv}$ being a group homomorphism means that
$\mathrm{Inv}(f\circ g)=\mathrm{Inv}(f)+\mathrm{Inv}(g)$
for any two interval exchange transformations $f$, $g$ of the same interval. It follows immediately that $\mathrm{Inv}(f^{n})=n\mathrm{Inv}(f)$, hence $f$ being periodic implies that $\mathrm{Inv}(f)=0$. Also, it follows that $\mathrm{Inv}(g\circ f\circ g^{-1})=\mathrm{Inv}(f)$, that is, $\mathrm{Inv}$ is a conjugacy invariant.
Recall that $\mathbb{R}\otimes_{\mathbb{Q}}\mathbb{R}$ is a vector space over $\mathbb{R}$ whose basis over $\mathbb{R}$ can be identified with a basis of $\mathbb{R}$ over $\mathbb{Q}$. More precisely: choosing any basis $\mathfrak{B}$ of $\mathbb{R}$ over $\mathbb{Q}$ yields an identification of $\mathbb{R}\otimes_{\mathbb{Q}}\mathbb{R}$ with the real vector space freely generated by $\mathfrak{B}$. Thus, $\mathbb{R}\otimes_{\mathbb{Q}}\mathbb{R}$ and $\mathbb{R}\wedge_{\mathbb{Q}}\mathbb{R}$ are infinite dimensional real vector spaces. Any invariant with values in these spaces can be alternatively viewed as an infinite collection of real-valued invariants. For example, choosing any $\mathbb{Q}$-bilinear map $\beta:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ yields the corresponding invariant
$\mathrm{Inv}_{\beta}(g):=\beta(\lambda_{1},\tau_{1})+\beta(\lambda_{2},\tau_{2})+\dots+\beta(\lambda_{n},\tau_{n}).$
It is straightforward to define interval exchange transformations of any metric oriented graph (a graph whose edges carry a Euclidean metric and an orientation). Also, since changing an interval exchange transformation at finitely many points does not change the corresponding element of the group $\mathrm{IET}(I)$, it is also irrelevant what the domain of a given interval exchange map looks like as long as it is a finite metric graph. For this reason, we will speak freely of interval exchange maps defined on various graphs.
Next, we want to generalize the SAF invariant to higher dimensional polytope exchange transformations. Specifically, define a dynamic invariant as a group homomorphism from $\mathrm{PET}(P)$ to $\mathbb{R}$. In what follows, new dynamic invariants will be introduced in 2D (they also make sense in dimensions $>2$).
2.5. The ideal Julia set
In this section, $f$ stands for a polygon exchange transformation of $\Delta$.
Definition 2.19 (Essential nest).
A nest $(Q_{n})$ is said to be essential if $Q_{n}$ is an essential set (see Definition 2.6), for every $n\geqslant 0$. Say that $(Q_{n})$ is periodic if $Q_{n+1}=Q_{n}$ for all sufficiently large $n$; otherwise $(Q_{n})$ is said to be non-periodic.
It turns out that essential and non-periodic are equivalent notions.
Lemma 2.20.
A nest $(Q_{n})$ is essential if and only if it is non-periodic.
Proof.
If $(Q_{n})$ is a periodic nest, then $Q_{n}$ is a periodic domain, for large $n$, hence not essential. Conversely, if $Q_{n}$ is not essential, then all non-boundary points of $Q_{n}$ are periodic, and there is a positive integer $m\geqslant n$ such that the period of any periodic point in $Q_{n}$ divides $m$. We claim that $Q_{m+1}=Q_{m}$. Otherwise, there is a point of $\Gamma_{m+1}$ inside $Q_{m}$. Nearby points must map close to $\Gamma_{1}$ under $f^{m}$. But $f^{m}$ sends on all non-boundary points of $Q_{m}\subset Q_{n}$ to themselves, hence the only points near $\Gamma_{m+1}$ in $Q_{m}$ are those near $\Gamma_{1}$, in particular, near the boundary of $Q_{m}$. ∎
Essential nests give another characterization of the Julia set.
Lemma 2.21.
A point $z\in\overline{\Delta}$ lies in the impression of an essential nest if and only if $z\in J$. Moreover, any essential level $n>0$ bi-piece belongs to some essential nest.
Proof.
Evidently, if $z$ is in the impression of an essential nest, then $z\in J$. We need only prove the converse.
Pick a Julia point $z$. An open set $O$ is said to be essential near $z$ if the intersection of an arbirtary neighborhood of $z$ with $O$ is essential. Clearly, if $O$ is essential near $z$, then $z\in\overline{O}$ even though $z$ may not lie in $O$. Polygons $Q_{n}$ can be defined inductively as follows. Set $Q_{0}=\Delta$ and, for $n>0$, define $Q_{n}$ as a level $n$ bi-piece that is contained in $Q_{n-1}$, and that is essential near $z$. The existence of such $Q_{n}$ follows from $Q_{n-1}$ being essential near $z$ and from the fact that there are only finitely many level $n$ bi-pieces in $Q_{n-1}$.
Let us now prove the last claim of the lemma. Take an essential level $n>0$ bi-piece $Q_{n}$. For $k\leqslant n$, define $Q_{k}$ as the level $k$ bi-piece containing $Q_{n}$. By induction, define $Q_{k}$ for $k>n$ as any essential level $k$ bi-piece contained in $Q_{k-1}$. Since $Q_{k-1}$ is essential (by induction), one can choose an essential level $k$ piece contained in $Q_{k-1}$. In this way, one constructs an essential nest such that $Q_{n}$ belongs to it. ∎
Nests form a compact metric space, on which a natural homemorphism acts induced by $f$.
Definition 2.22 (The ideal Julia set).
Define the ideal Julia set $\mathcal{J}=\mathcal{J}(f)$ of $f$ as the set of all essential nests. This is not only a set but a metric space, in which the distance between two nests $(Q_{n})$ and $(Q^{\prime}_{n})$ is by definition $2^{-m}$, where $m$ is the maximal non-negative integer such that $Q_{m}=Q^{\prime}_{m}$ (it follows that $Q_{n}=Q^{\prime}_{n}$ for all $n\leqslant m$). It is not hard to see that this distance function is indeed a metric, even an ultra-metric: given three nests, the distance from the first to the third is at most the maximum of the distance from the first to the second and the distance from the second to the third.
A basis of open sets in $\mathcal{J}$ is labeled by essential bi-pieces $Q$ of arbitrary level $n$. Given such $Q$, the corresponding basis open set $\mathcal{B}(Q)$ is formed by all essential nests $(Q_{k})$ with $Q_{n}=Q$.
Lemma 2.23.
The ideal Julia set $\mathcal{J}$ is compact.
Proof.
It suffices to show that any cover of $\mathcal{J}$ by basis open sets admits a finite subcover. By way of contradiction, consider a family $(\mathcal{B}(Q^{\alpha}))$ of basis open sets covering the space $\mathcal{J}$ and such that it has no finite subcover. Here $\alpha$ ranges through some index set. Now define an essential nest $(Q_{n})$ as follows. Set $Q_{0}=\Delta$ and, for every $n>0$, define $Q_{n}$ as a level $n$ essential bi-piece contained in $Q_{n-1}$ and such that the given cover has no finite subcover over $\mathcal{B}(Q_{n})$. Finding such $Q_{n}$ is possible since $\mathcal{B}(Q_{n-1})$ is a finite union of $\mathcal{B}(Q)$, with $Q$ ranging through all level $n$ bi-pieces contained in $Q_{n-1}$. Evidently, the thus obtained nest $(Q_{n})$ is essential, since each $Q_{n}$ is essential (otherwise, $\mathcal{B}(Q_{n})$ would be empty, hence admitting a finite subcover). A contradiction with the fact that is covered by some $\mathcal{B}(Q^{\alpha})$, that is, $Q_{n}=Q^{\alpha}$ for some $n$: then $\mathcal{B}(Q_{n})$ is covered by a single basis open set from the given cover, contrary to the choice of $Q_{n}$. ∎
The polygon exchange transformation $f$ gives rise to a homeomorphism $h_{f}:\mathcal{J}\to\mathcal{J}$ as follows. By definition, the image of an essential nest $(Q_{n})$ under $h_{f}$ is the essential nest $(Q^{\prime}_{n})$, where $Q^{\prime}_{n}$ is the level $n$ bi-piece containing $f(Q_{n+1})$. It is easy to see that $h_{f}$ is indeed a homeomorphism by looking at how $h_{f}$ acts on the basis open sets $\mathcal{B}(Q)$. Results of this section are summarized below.
Theorem 2.24.
Suppose that $f$ is a non-periodic polygon exchange transformation. The ideal Julia set $\mathcal{J}$ of $f$ is compact and nonempty. Moreover, $h_{f}$ is a self-homeomorphism of $\mathcal{J}$ without periodic points.
Proof.
Firstly, $\mathcal{J}$ is compact by Lemma 2.23, and is nonempty by Lemmas 2.8, 2.21. Let us now show that $h_{f}$ has no periodic points. A periodic point for $h_{f}$ is an essential nest $(Q_{n})$ with the property that $f^{k}(Q_{n+k})\subset Q_{n}$ for a fixed $k>0$ and all $n>0$. Denote by $g$ the translation by which $f^{k}$ acts on $Q_{n+k}$ for all sufficiently large $n$. If $g$ is the identity, then the nest $(Q_{n})$ is periodic rather than essential. Assume now that $g$ is not the identity. Choose an integer $m>0$ so large that $g^{m}(\Delta)\cap\Delta=\varnothing$. This is a contradiction with the property $f^{n+mk}(Q_{n+mk})\subset Q_{n}$. Hence, $h_{f}$ acts without periodic points. ∎
All results about the ideal Julia set can be rephrased in terms of symbolic dynamics. Consider the alphabet $\mathcal{A}$ whose symbols are in one-to-one correspondence with components of $\mathrm{dom}(f)$; one can formally set $\mathcal{A}:=\pi_{0}(\mathrm{dom}(f))$. Every nest $(Q_{n})$ gives rise to a bi-infinite sequence in the alphabet $\mathcal{A}$, namely, the $n$-th symbol is the component of $\mathrm{dom}(f)$ containing $f^{n-1}(Q_{n})$ if $n>0$ and the component of $\mathrm{dom}(f)$ containing $f^{n-1}(Q_{-n+1})$ if $n\leqslant 0$. This correspondence embeds the dynamical system $(\mathcal{J},h_{f})$ into $(\mathcal{A}^{\mathbb{Z}},\sigma)$, where $\sigma$ is the shift map.
Remark 2.25 (No long repetitions).
The fact that $\mathcal{J}$ is compact, and $h_{f}$ acts on $\mathcal{J}$ without periodic points, is translated into symbolic dynamics language as follows. A symbolic sequence from $\mathcal{A}^{\mathbb{Z}}$ representing a given element of $\mathcal{J}$ cannot have subwords that are arbitrarily long repetitions of the same word. More precisely, for any given finite word $\mathsf{w}$ in the alphabet $\mathcal{A}$, there is an integer $n>0$ with the property that at most $n$ copies of $\mathsf{w}$ can appear next to each other in a symbolic sequence from $\mathcal{A}^{\mathbb{Z}}$ representing an element of $\mathcal{J}$.
2.6. Valuations
We first recall some classical notions and results related to valuations in a non-dynamical setting. The following definition is standard.
Definition 2.26 (Valuations).
Let $\mathcal{P}$ be the set of all compact convex polygons in the plane, including $\varnothing$. A valuation on convex polygons with values in $\mathbb{R}$ is by definition a function $\mu:\mathcal{P}\to\mathbb{R}$ such that $\mu(\varnothing)=0$ and
$\mu(P\cup Q)=\mu(P)+\mu(Q)-\mu(P\cap Q)$
for convex polygons $P$, $Q$ such that $P\cup Q$ is also convex. This property of valuations is called finite additivity. A valuation is therefore the same as a finitely additive (signed) measure.
It is straightforward to check that any valuation $\mu$ extends to arbitrary, not necessarily convex, bounded polygons, and satisfies the same finite additivity property on this larger class of sets.
A valuation $\mu$ is said to be simple if $\mu(P)=0$ whenever $\dim(P)<2$. Say that $\mu$ is translation invariant if $\mu(P+\vec{v})=\mu(P)$ for all convex polygons $P$ and all $\vec{v}\in\mathbb{R}^{d}$. Here, $P+\vec{v}$ means the result of the translation of $P$ by the vector $\vec{v}$. More generally, one can speak of $\mathsf{G}$-invariant valuations, where $\mathsf{G}$ is any group of affine transformations of $\mathbb{R}^{d}$. Simple $\mathsf{G}$-invariant valuations are useful for solving $\mathsf{G}$-scissors congruence problems. Namely, if $P$ and $Q$ are $\mathsf{G}$-scissors congruent, then any simple $\mathsf{G}$-invariant valuation takes the same value on $P$ and $Q$. In other words, simple invariant valuations help distinguish pairs of polytopes that are not scissors congruent.
Example 2.27 (Area).
Fix an area form on $\mathbb{R}^{2}$. Then one can assign the area $\mathrm{area}(P)$ to any polytope $P$. Area is a simple valuation invariant under the full group of isometries of $\mathbb{R}^{2}$. Now let $\xi:\mathbb{R}\to\mathbb{R}$ be any $\mathbb{Q}$-linear function. Then $\xi\circ\mathrm{area}$ is also a simple isometry invariant valuation; this construction gives uncoutably many such valuations, linearly independent over $\mathbb{R}$. Perimeter (suitably extended to degenerate polygons) is also an isometry invariant valuation, however, it is not simple.
Example 2.28 (Hadwiger–Glur valuations).
Fix a line $L\subset\mathbb{R}^{2}$ together with a co-orientation, that is, a labeling of the two components of $\mathbb{R}^{2}\setminus L$ as the positive and the negative sides of $L$. Given a co-orientation of $L$, we can consistently define the positive and the negative sides of any line parallel to $L$. Let now $P$ be a bounded polygon, not necessarily convex, and define $\mu(P)$ as follows. Every edge $e$ of $P$ parallel to $L$ contributes an additive term ±$\pm|e|$ to $\mu(P)$, where $|e|$ is the length of $e$, and the sign is plus or minus depending on whether $P$ is on the positive or on the negative side of $e$ (locally near $e$). The sum $\mu(P)$ of such terms is called the Hadwiger invariant of $P$, and the valuatuon $\mu:P\mapsto\mu(P)$ is called the Hadwiger–Glur valuation. It was introduced by Hadwiger and Glur in [HG51], see also [Had13]. Observe that $\mu$ is simple and translation invariant. Two polygons are $\mathsf{T}$-scissors congruent if and only if they have the same area and the same value of the Hadwiger invariants for all $L$, cf. [HG51]. Even though there are uncountably many Hadwiger–Glur valuations, only finitely many of them take nonzero values at a given polygon.
Let $P$ be a bounded polygon. By a dynamic invariant, we mean a group homomorphism $\Psi$ from $\mathrm{PET}(P)$ to the additive group of some $\mathbb{Q}$-vector space. For any two transformations $f$, $g\in\mathrm{PET}(P)$, we therefore have
$\Psi(f\circ g)=\Psi(f)+\Psi(g).$
It follows that $\Psi(f^{n})=n\Psi(f)$ for every $n\in\mathbb{Z}$. Therefore, if $f$ is periodic, then any dynamic invariant must vanish on $f$. Dynamic invariants are, indeed, congugacy invariants ($\Psi(g\circ f\circ g^{-1})=\Psi(f)$). Below, we give a general construction of multidimensional dynamic invariants.
Suppose that $\mu$ is a simple translation invariant valuation on polygons in $\mathbb{R}^{2}$. Take any $f\in\mathrm{PET}(P)$, whose domain consists of open convex polygons $P_{i}$ (where $i$ runs from $1$ to $n$) and whose restriction to $P_{i}$ is the translation by a vector $\vec{v}_{i}$. Define
$\mathrm{Inv}_{\mu}(f)=\sum_{i=1}^{n}\mu(P_{i})\otimes\vec{v}_{i}\in\mathbb{R}\otimes_{\mathbb{Q}}\mathbb{R}^{2}$
and call it the $\mu$-invariant of $f$. The main property of this invariant is the following.
Theorem 2.29.
The function $\mathrm{Inv}_{\mu}$ is a group homomorphism:
$\mathrm{Inv}_{\mu}(g\circ f)=\mathrm{Inv}_{\mu}(g)+\mathrm{Inv}_{\mu}(f).$
In particular, any $\mathbb{Q}$-bilinear functional $\beta:\mathbb{R}\times\mathbb{R}^{2}\to\mathbb{C}$ defines a dynamic invariant $\mathrm{Inv}_{\mu,\beta}$ with values in $\mathbb{C}$, which can be computed as the sum of $\beta(\mu(P_{i}),\vec{v}_{i})$ for $i=1$, $\dots$, $n$.
Proof.
Suppose that the domain of $f$ has components $P_{1}$, $\dots$, $P_{n}$, and the domain $g$ has components $Q_{1}$, $\dots$, $Q_{m}$. Write $f_{i}:=f|_{P_{i}}$ and $g_{j}:=g|_{Q_{j}}$. Note that the components of $\mathrm{dom}(g\circ f)$ are those of the sets $P_{ij}=f_{i}^{-1}(Q_{j})$ which are nonempty. Let $\vec{v}_{i}$, $\vec{w}_{j}$ stand for the translation vectors of $f_{i}$, $g_{j}$, respectively, and $\mathfrak{S}$ for the set of pairs $(i,j)$ such that $P_{ij}\neq\varnothing$. Also, for every $j=1$, $\dots$, $m$, let $\mathfrak{S}_{j}$ stand for the set of $i$s with $(i,j)\in\mathfrak{S}$, and write $\mathfrak{S}^{i}$ for the set of all $j$ with $(i,j)\in\mathfrak{S}$. Then, on $P_{ij}$, the composition $g\circ f$ acts as a translation by vector $\vec{v}_{i}+\vec{w}_{j}$. Computing $\mathrm{Inv}_{\mu}(g\circ f)$ by definition yields
$\sum_{(i,j)\in\mathfrak{S}}\mu(P_{ij})\otimes(\vec{v}_{i}+\vec{w}_{j})=\sum_{i=1}^{n}\sum_{j\in\mathfrak{S}^{i}}\mu(P_{ij})\otimes\vec{v}_{i}+\sum_{j=1}^{m}\sum_{i\in\mathfrak{S}_{j}}\mu(P_{ij})\otimes\vec{w}_{j}.$
Fix $j$ and consider each term of the right hand side for this fixed value of $j$. Note that the sum of $\mu(P_{ij})$ over all $i\in\mathfrak{S}_{j}$ equals $\mu(Q_{j})$, by translation invariance and simplicity of the valuation $\mu$. It follows that the second term in the right hand side equals the sum of $\mu(Q_{j})\otimes\vec{w}_{j}$ over all $j$, which is $\mathrm{Inv}_{\mu}(g)$ by definition of ${\mu}$-invariant. Similarly, the sum of $\mu(P_{ij})$ over all $j\in\mathfrak{S}^{i}$ equals $\mu(P_{i})$, by translation invariance and simplicity of $\mu$. The first term in the right hand side is then the sum over all $i$ from $1$ to $n$ of $\mu(P_{i})\otimes\vec{v}_{i}$, which is $\mathrm{Inv}_{\mu}(f)$ by definition of ${\mu}$-invariant. The theorem follows. ∎
The following corollary of Theorem 2.29 will be used in the proof of Theorem 1.1.
Corollary 2.30.
Let $\Delta$ be a bounded polygon, and take $f\in\mathrm{PET}(\Delta)$. If $\mathrm{Inv}_{\mu}(f)\neq 0$ for some simple $\mathsf{T}$-invariant valuation $\mu$, then $f$ has infinite order in the group $\mathrm{PET}(\Delta)$.
Proof.
Similarly to the case of interval exchange maps, the relation $f^{n}=id$ in the group $\mathrm{PET}(\Delta)$ implies $n\,\mathrm{Inv}_{\mu}(f)=0$ by Theorem 2.29, hence $\mathrm{Inv}_{\mu}(f)=0$, a contradiction. ∎
Next, we discuss dynamical version of the Hadwiger–Glur valuation.
Definition 2.31 (Dynamic Hadwiger invariant).
Fix a co-oriented line $L$ in $\mathbb{R}^{2}$, and define the dynamic Hadwiger invariant $\mathrm{Inv}_{L}$ associated with $L$ as $\mathrm{Inv}_{\mu_{L}}$, where $\mu_{L}$ is the Hadwiger–Glur valuation corresponding to $L$. Recall that $\mu_{L}$ takes values in $\mathbb{R}\otimes_{\mathbb{Q}}\mathbb{R}^{2}$.
For a given polygon exchange transformation $f$, the dynamic Hadwiger invariant $\mathrm{Inv}_{L}(f)$ associated with $L$ can be computed as follows. A component $Q$ of $\mathrm{dom}(f)$ and an edge $E$ of $Q$ contribute the additive term $[E|Q]:=\mathrm{sign}[E|Q](|E|\otimes\vec{v})$ to the invariant, where $E$ is an edge of $Q$ parallel to $L$ (we write $|E|$ for the length of $E$), and $\vec{v}$ is the translation vector of $f|_{Q}$. The sign $\mathrm{sign}[E|Q]$ is defined as $+1$ if $Q$ lies on the positive side of $E$ and as $-1$ otherwise. A pair $(Q,E)$ like that, i.e., with $Q$ being a component of $\mathrm{dom}(f)$ and $E$ being an edge of $Q$ parallel to $L$, is called an $L$-adapted pair. It is said to be positive or negative according to whether $\mathrm{sign}[E|Q]$ is positive or negative. The following lemma summarizes the above discussion.
Lemma 2.32.
If $f$ is a polygon exchange transformation, then $\mathrm{Inv}_{L}(f)$ can be computed as the sum of $[E|Q]$ over all $L$-adapted pairs $(Q,E)$.
To obtain a non-trivial invariant, it is necessary to consider either the tensor product over the rationals, or $\mathbb{Q}$-bilinear functions that are not linear over $\mathbb{R}$, as shown by the following proposition (due to Victor Kleptsyn, private communication).
Proposition 2.33.
The dynamic Hadwiger invariant lies in the kernel of the multiplication map $\mathfrak{m}:\mathbb{R}\otimes_{\mathbb{Q}}\mathbb{R}^{2}\to\mathbb{R}^{2}$ given by $\mathfrak{m}(x\otimes\vec{v})=x\vec{v}$.
Proof.
Let $f$ be a polygon exchange transformation of a bounded polygon $P$ so that $\mathrm{dom}(f)$ consists of open polygons $P_{1}$, $\dots$, $P_{n}$, and $f|_{P_{i}}$ is the translation by vector $\vec{v}_{i}$. Observe that $\mathfrak{M}(f):=\mathfrak{m}(\mathrm{Inv}_{L}(f))$ can be computed as the sum of $\mu_{L}(P_{i})\vec{v}_{i}$. We want to show that $\mathfrak{M}(f)=0$.
Given a polygon $Q$, let $\mathfrak{I}(Q)$ be the integral over the boundary $\partial Q$ of $Q$ of the vector-valued function equal to ±$\pm z$ at points $z\in\partial Q$ such that $z$ lies in a support line of $Q$ parallel to $L$ (the sign is plus or minus depending on whether $Q$ lies on the positive or on the negative side of the support line) and equal to 00 at all other points. Integration is performed against the usual Euclidean length measure on $\partial Q$. Note that, for every vector $\vec{v}\in\mathbb{R}^{2}$,
$\mathfrak{I}(Q+\vec{v})-\mathfrak{I}(Q)=\mu_{L}(Q)\vec{v},$
as follows directly from the definition of the Hadwiger invariant $\mu_{L}(Q)$. Clearly, $\mathfrak{I}$ is additive with respect to subdivision: if $Q$ is cut into pieces $Q_{1}$, $\dots$, $Q_{k}$, then $\mathfrak{I}(Q)$ equals the sum of $\mathfrak{I}(Q_{j})$ for $j=1$, $\dots$, $k$. (Every boundary interval of a piece $Q_{j}$ that is shared with another piece $Q_{j^{\prime}}$ contributes two terms to $\mathfrak{I}(Q)$, which differ only by a sign and therefore cancel each other).
It now follows that
$\mathfrak{I}(P)=\sum_{i=1}^{n}\mathfrak{I}(P_{i}+\vec{v}_{i})=\sum_{i=1}^{n}\mathfrak{I}(P_{i})+\mu_{L}(P_{i})\vec{v}_{i}=\mathfrak{I}(P)+\mathfrak{M}(f),$
which implies that $\mathfrak{M}(f)=0$. Here, we used the additivity of $\mathfrak{I}$ two times. ∎
3. Bounded restrictions of outer billiards
Write $\Pi=\Pi_{N}$ for the regular $N$-gon inscribed in the unit circle such that one of its vertices lies in the positive $x$-semiaxis. Under a suitable identification of $\mathbb{R}^{2}$ with $\mathbb{C}$ (so that the origin of $\mathbb{R}^{2}$ identifies with the zero complex number, and the positive semiaxis identifies with the set of positive real numbers), the vertices of $\Pi$ are precisely all $N$-th roots of unity; they have the form $\zeta^{k}$ for $\zeta:=\exp(2\pi\mathbf{i}/N)$ and $k=0$, $\dots$, $N-1$. Recall a general result of Rukhovich [Ruk22] (see also the “Twice-odd Lemma” of [Hug13] for an earlier experimental evidence).
Theorem 3.1 ([Ruk22]).
Let $N>3$ be odd. The map $f_{\Pi_{N}}$ has aperiodic points if and only if $f_{\Pi_{2N}}$ does.
Thus, it suffices to prove Theorem 1.1 only for even values of $N$. For this reason, we assume from now on and throughout the rest of the paper that $N$ is even.
3.1. Vassal polygons
Section 3.1 claims no originality. In “setting the stage” below, we essentially follow the construction of [GS92, Tab93, BC11, Ruk22], however, the notation and some conventions will be different. Let $R$ be the counterclockwise rotation by $\theta_{N}:=2\pi/N$ about 0, that is, $R(z):=\zeta z$ for all $z\in\mathbb{C}$. Recall that $V_{0}$ is a sector with vertex $1=\zeta^{0}$ formed by lines containing the two incident to $1$ edges of $\Pi$, see Fig. 1. Then $V_{k}=R^{k}(V_{0})$; here the index $k$ can be viewed as a residue modulo $N$. Write $S_{k}$ for the central reflection (=half-turn) about the vertex $\zeta^{k}$; we have
$S_{k}(z)=2\zeta^{k}-z=R^{k}\circ S_{0}\circ R^{-k}(z).$
Consider the strip formed by the lines $1\zeta$ and $(-1)(-\zeta)$; call it strip 0. The line $(-\zeta)(-\zeta^{2})$ divides strip 0 into two equal unbounded parts. One part, call it the senior part, contains the polygon $\Pi$. Define the vassal part as the other part of strip 0; it is centrally symmetric to the senior part with respect to the midpoint of the common edge. A regular polygon $\Pi^{\dagger}$ in the vassal part is symmetric to $\Pi$; call it the vassal polygon of $\Pi$. Since the two polygons $\Pi$ and $\Pi^{\dagger}$ are symmetric to each other with respect to the midpoint of the common boundary between the senior and the vassal parts of strip 0, these polygons have the same sidelength $\lambda_{0}=|\zeta-1|$.
Recall that $N$ is even, hence it has the form $N=2m+2$. Let $c_{m}$ be the intersection point of the line $(-\zeta)(-\zeta^{2})$ with the line $1\zeta$. Polygons $\Pi$ and $\Pi^{\dagger}$ are symmetric with respect to the midpoint of the segment $[-\zeta;c_{m}]$. It follows that the line $(-\zeta)(-\zeta^{2})$ contains an edge $[c_{m-1};c_{m}]$ of $\Pi^{\dagger}$. Write $c_{0}$, $\dots$, $c_{m}$ for the vertices of $\Pi^{\dagger}$ symmetric to $1$, $\zeta^{-1}$, $\dots$, $\zeta^{-m}=-\zeta$, so that $c_{k}$ is symmetric to $\zeta^{-k}$. Both $\Pi$ and $\Pi^{\dagger}$ are symmetric with respect to the reflection axis of strip 0. Hence the line $1\zeta^{-1}$ contains the edge $[c_{0};c_{1}]$ of $\Pi^{\dagger}$.
Lemma 3.2.
The polygons $\Pi^{\dagger}$ and $R^{-1}\Pi^{\dagger}$ have $c_{0}$ as a common vertex.
Proof.
See Fig. 2, left. Define strip $-1$ as the image of strip 0 under $R^{-1}$. It is bounded by the lines $\zeta^{-1}1$ and $(-1)(-\zeta)$. Note that $R^{-1}c_{m}=c_{0}$ since $c_{0}$ is on the boundary line $\zeta^{-1}1$ of strip $-1$ corresponding to the boundary line $1\zeta$ of strip 0 under $R^{-1}$ and, moreover, the distance from $c_{0}$ to $\zeta^{-1}$ equals the distance from $c_{m}$ to 1. It follows that the vertex $c_{0}$ of $\Pi^{\dagger}$ is also a vertex of the polygon $R^{-1}\Pi^{\dagger}$, as claimed. ∎
Figure 2. Left: the vassal polygon $\Pi^{\dagger}$, strip 0 and strip $-1$. Right: the polygon $\Pi$ and the piece $A_{0}$ for $N=8$.
It is easy to characterize the vassal polygon in terms of dynamics: the following lemma is straightforward and is left to the reader.
Lemma 3.3.
The interior of the vassal polygon $\Pi^{\dagger}$ of $\Pi$ coincides with the set of all points $z\in V_{0}$ with the property that $f_{\Pi}^{i}(z)\in V_{mi}$ for all integer values of $i$, where the subscript $mi$ of $V_{mi}$ is understood as a residue modulo $N$.
The $R$-orbit of $\Pi^{\dagger}$ consists of $N$ regular polygons touching each other in a cyclic order. Hence the union of the closures of these polygons defines a bounded complementary domain $X$ containing $\Pi$. Set $A_{0}:=V_{0}\setminus\overline{\Pi}^{\dagger}$; this is a non-convex polygon with vertices $1$, $c_{1}$, $\dots$, $c_{m}$. Here, the numbering of the vertices follows their cyclic order in the boundary of $A_{0}$, see Fig. 2, right.
Note that any diameter line of $\Pi$ is a symmetry axis of $X$, and that $f_{\Pi}$ takes $A_{0}$ to the symmetric polygon with respect to the diameter through $1$ and $-1$. In particular, $f_{\Pi}(A_{0})\subset X$, and similarly for all pieces $A_{k}:=R^{k}A_{0}$ for $k=0$, $\dots$ $N-1$. The restriction of $f_{\Pi}$ to $X\setminus\Pi$ is therefore a piecewise rotation of $X$ with pieces $A_{k}$, cf. [Ruk22, Lemma 20]. Denote it by $f_{X}$.
Let $M$ denote the half-turn about the origin, that is, the map $z\mapsto-z$; one can also write $M=R^{N/2}$. Since $f_{X}$ acts as a half-turn on every component of $\mathrm{dom}(f_{X})$, the map $g_{X}:=M\circ f_{X}$ is clearly a polygon exchange transformation. Fig. 3 illustrates the action of $g_{X}$. Observe that $M$ conjugates $f_{X}$ with itself, therefore, it takes orbits to orbits, hence $g_{X}$ has aperiodic points if and only if $f_{X}$ does.
Figure 3. The action of the map $g=g_{X}$ on $X$. Left: components of $\mathrm{dom}(g)$. Right: their $g$-images.
3.2. Computation of $\mathrm{Inv}_{L}(g_{X})$
All tensor products we consider are over $\mathbb{Q}$; for this reason, we will from now on omit $\mathbb{Q}$ from the notation. For our subsequent computation of the dynamic Hadwiger invariant for $g_{X}$, we will need to know the lengths of all edges of $A_{0}$. Edges $[c_{1};c_{2}]$, $\dots$, $[c_{m-1};c_{m}]$ all have the same length as an edge of $\Pi$. Namely, their length is $\lambda_{0}=2\sin\frac{\theta}{2}$. Side edges $[1,c_{1}]$ and $[1;c_{m}]$ of $A_{0}$ have the same length $\lambda^{\prime}$, which we will now compute. Consider the triangle with vertices $1$, $-1$, $c_{0}$; this is an isosceles triangle with base length 2 and angles $\theta$, $\frac{\pi-\theta}{2}$, $\frac{\pi-\theta}{2}$. It follows that the side edge of this triangle has length
$\lambda_{0}+\lambda^{\prime}=\frac{1}{\sin\frac{\theta}{2}}=\frac{2}{\lambda_{0}},$
which gives an expression for $\lambda^{\prime}$ in terms of $\lambda_{0}$. In the sequel, we will need an expression for $\lambda^{\prime}$ in terms of $\lambda_{0}$ and $\zeta$. One can use the relation $\zeta+\zeta^{-1}=2-\lambda_{0}^{2}$, in which the left-hand side is equal to $2\cos\theta$, and the right-hand side follows from the well-known expression for $\cos\theta$ in terms of $\sin\frac{\theta}{2}$. As a consequence, we obtain that
$\lambda^{\prime}=\lambda_{0}\frac{\zeta+\zeta^{-1}}{2-\zeta-\zeta^{-1}}.$
A dynamic Hadwiger invariant of $g_{X}$ is associated with some choice of a line $L$. Choose the line $L$ through the vertices $1=\zeta^{0}$ and $\zeta^{1}$ of $\Pi$, and define the positive side of $L$ as the one not containing $\Pi$. Each edge $e$ of $A_{0}$ gives rise to two $L$-adapted pairs (see Section 2.6) $(A_{k},E)$ and $(M(A_{k}),M(E))$, where $E=R^{k}e$ is the corresponding to $e$ edge of $A_{k}$ parallel to $L$, and $M$ is the map $z\to-z$. Call the pairs $(A_{k},E)$ and $(MA_{k},ME)$ the $L$-adapted pairs corresponding to $e$. Among these two pairs, exactly one is positive, and the other one is negative. For example, $(A_{0},[c_{m};1])$ is a negative pair, hence, the corresponding to the edge $[c_{m};1]$ positive pair is $(A_{m+1},M[c_{m},1])$. Positive pairs are described in the next lemma.
Lemma 3.4.
All positive $L$-adapted pairs are
$(A_{1},R[1,c_{1}]),\quad(A_{m+1},M[c_{m},1]),$
$(A_{-k},R^{-k}[c_{m-k},c_{m-k+1}]),\quad k=1,\dots,m-1.$
Proof.
In order to obtain a segment parallel to $L$, rotate $[c_{m-1};c_{m}]$ by $R^{-1}$. Clearly, $A_{-1}$ appears on the positive side of $R^{-1}[c_{m-1};c_{m}]$. Next, each time we pass from $k$ to $k+1$, the next segment $[c_{m-k-1};c_{m-k}]$ is rotated by angle $-\theta$ compared to $[c_{m-k};c_{m-k+1}]$, and the side on which $A_{0}$ lies rotates accordingly. This implies the desired. ∎
We also need to know the translation vectors of $g_{X}$.
Lemma 3.5.
Each piece $A_{k}$ is translated by $\vec{v}_{k}=-2\zeta^{k}$ under $g_{X}$.
Proof.
Under the map $f_{X}$, or, more precisely, the continuous extension of it to $\overline{A}_{k}$, the vertex $\zeta^{k}$ of $V_{k}$ is mapped to itself. Next, $M$ takes $\zeta^{k}$ to $M\zeta^{k}=-\zeta^{k}$. Therefore, the continuous extension of $g_{X}|_{A_{k}}$ onto $\overline{A}_{k}$ takes $\zeta^{k}$ to $-\zeta^{k}$. In other words, $g_{X}$ translates the piece $A_{k}$ by vector $\vec{v}_{k}=(-\zeta^{k})-\zeta^{k}=-2\zeta^{k}$. ∎
Now that all required edge lengths and translation vectors are found, we can complete the computation of the dynamic Hadwiger invariant.
Theorem 3.6.
Suppose $N>2$ is even. The dynamic Hadwiger invariant of $g_{X}$ equals
$\mathrm{Inv}_{L}(g_{X})=4\lambda_{0}\otimes\frac{\zeta(\zeta+\zeta^{-1})}{1-\zeta}-4\lambda_{0}\frac{\zeta(\zeta+\zeta^{-1})}{(1-\zeta)^{2}}\otimes(1-\zeta).$
In the expressions for $\mathrm{Inv}_{L}(g_{X})$ given above, we used the following convention: if $x$ and $z$ are products of complex numbers, then these products are evaluated first, before evaluating $x\otimes z$. The same convention is used throughout the paper.
Proof.
Consider an edge of $A_{0}$ and the two corresponding $L$-adapted pairs. Observe that the corresponding pieces $A_{k}$ and $MA_{k}$ are translated by opposite vectors under $g_{X}$, by Lemma 3.5. For this reason, $\mathrm{Inv}_{L}(g_{X})$ is twice the sum of the contributions of all positive pairs. Edges of $A_{0}$ not on the boundary of $V_{0}$ contribute the sum $2\lambda_{0}\otimes\vec{v}_{\mathrm{bot}}$, where $\vec{v}_{\mathrm{bot}}$ is the sum of $\vec{v}_{-k}$ over all $k=1$, $\dots$, $m-1$, and the subscript ‘bot’ stands for ‘bottom’.
Computing $\vec{v}_{\mathrm{bot}}$ amounts to summing a geometric series, namely,
$\vec{v}_{\mathrm{bot}}=-2\sum_{k=1}^{m-1}\zeta^{-k}=-2\zeta^{-m}\frac{\zeta-\zeta^{m}}{1-\zeta}=2\frac{\zeta(\zeta+\zeta^{-1})}{1-\zeta}.$
It remains to account for the contributions of the two edges of $A_{0}$ that lie on the boundary of $V_{0}$. Write this last contribution as $2\lambda^{\prime}\otimes\vec{v}_{\mathrm{top}}$; where $\vec{v}_{\mathrm{top}}$ is equal to $\vec{v}_{0}+\vec{v}_{1}=2-2\zeta$. Using the expression for $\lambda^{\prime}$ through $\lambda_{0}$, we obtain the desired expression for $\mathrm{Inv}_{L}(g_{X})$. ∎
It remains to verify that the invariant is non-zero.
Corollary 3.7.
If $N>3$ is even, then $\mathrm{Inv}_{L}(g_{X})\neq 0$.
Proof.
According to the proof of Theorem 3.6, the invariant $\mathrm{Inv}_{L}(g_{X})$ can be written as $2\lambda_{0}\otimes\vec{v}_{\mathrm{bot}}+2\lambda^{\prime}\otimes\vec{v}_{\mathrm{top}}$. Here $\vec{v}_{\mathrm{bot}}$ and $\vec{v}_{\mathrm{top}}$ are complex numbers such that
$-\frac{\vec{v}_{\mathrm{bot}}}{\vec{v}_{\mathrm{top}}}=\frac{\lambda^{\prime}}{\lambda_{0}}=\frac{\cos\theta}{2-\cos\theta}.$
This is a real but not rational number, under our assumptions on $N$. Hence, $\vec{v}_{\mathrm{bot}}$ and $\vec{v}_{\mathrm{top}}$ are linearly independent over $\mathbb{Q}$. Let $\mathcal{V}$ be the rational vector space spanned by these two vectors. Note that $\mathrm{Inv}_{L}(g_{X})$ lies in $\mathbb{R}\otimes\mathcal{V}$, and the latter group has a natural structure of a real vector space with basis $1\otimes\vec{v}_{\mathrm{bot}}$ and $1\otimes\vec{v}_{\mathrm{top}}$. Since $\mathrm{Inv}_{L}(g_{X})$ is a non-trivial linear combination of the basis vectors, this element is nonzero in $\mathbb{R}\otimes\mathcal{V}\subset\mathbb{R}\otimes\mathbb{C}$. ∎
By Theorem 2.29, a periodic polygon exchange transformation of a bounded polygon must have zero invariant $\mathrm{Inv}_{L}$. It now follows from Corollary 3.7 that $g_{X}$ is not periodic. Hence $f_{X}=M\circ g_{X}$ is not periodic either, since $M$ has period 2 and commutes with $f_{X}$ and $g_{X}$. We have thus proved the following theorem.
Theorem 3.8.
The scissors automorphism $f_{X}$ of $X$ is not periodic.
4. Boundary dynamics of piecewise rotations
Section 4 describes alternative methods for proving similar results to those of Section 3. While it can be useful for future research, this section is not directly involved in the proof of Theorem 1.1 and, for this reason, can be omitted. Given an orientation preserving scissors automorphism $f$ of a bounded polygon $P$, we will introduce an associated interval exchange transformation. Computing the SAF invariant of the latter transformation helps determining aperiodicity of the initial map $f$. Let $P_{1}$, $\dots$, $P_{m}$ be the components of $\mathrm{dom}(f)$. Define $\partial f$ as the disjoint union of the boundaries of all $P_{i}$. More precisely, elements of $\partial f$ are pairs of the form $(z,P_{i})$, where $z\in\partial P_{i}$. Then $\partial f$ is a finite graph that is a union of $m$ cycles. It has a natural orientation induced from the orientation of $\mathbb{R}^{2}$: every component of $\partial f$ corresponding to the boundary of some piece $P_{i}$ is oriented as the boundary, i.e., counterclockwise.
Note that $\partial f$ projects naturally onto $\overline{P}\setminus\mathrm{dom}(f)$; this projection is two-to-one except the points projecting to $\partial P$ and except finitely many vertices of $\partial f$. Every point of $\partial P$ that is not a vertex of a component of $\mathrm{dom}(f)$ has only one preimage in $\partial f$, and vertices of pieces $P_{i}$ may have more than two preimages.
4.1. Boundary quasi-commutators
Recall that, when dealing with interval exchange transformations, one ignores any effect happening only on a finite subset of points. In particular, it makes sense to consider the group $\mathrm{IET}(\partial f)$ of all interval exchange transformations of $\partial f$, even though $\partial f$ is not an interval (but one can make it into an interval by cutting and pasting at finitely many points). Note that $f$ defines a natural map $f^{*}:\partial f\to\partial f^{-1}$. More precisely, $f^{*}$ sends a point $(z,i)\in\partial f$ to the point $(\overline{f}_{i}(z),f(P_{i}))$, where $\overline{f}_{i}$ is the continuous extension of $f|_{P_{i}}$ to the closure of $P_{i}$. A disadvantage of $f^{*}$ is that this map cannot be iterated. However, below, we will define an iterable map based on $f^{*}$.
Definition 4.1 (A boundary involution).
Let $\partial_{in}f$ be the inner boundary of $f$, formed by (the copies of) the edges of $P_{i}$ not on the boundary of $P$. Call the remaining part of $\partial f$ the outer boundary and denote it by $\partial_{out}f$. Both $\partial_{in}f$ and $\partial_{out}f$ are viewed as subgraphs of $\partial f$. There is a natural involution $w:\partial_{in}f\to\partial_{in}f$ that takes $(z,i)$ to $(z,j)$ whenever $z$ is a common boundary point of $P_{i}$ and $P_{j}$. Informally speaking, $w$ changes sides. Note that $w$ is undefined at vertices of $P_{i}$ that lie inside $P$, however, this is immaterial for the definition of $w$ as a piecewise isometry of $\partial_{in}f$. Clearly, $w\circ w=id_{\partial_{in}f}$, except at finitely many points. Also, observe that $w$ reverses the orientation of $\partial_{in}f$. Extend $w$ to all of $\partial f$ as an involutive orientation reversing piecewise isometry. This is always possible: one just needs to define $w$ on $\partial P$ (except at finitely many points, $\partial P$ identifies with $\partial_{out}f$), for example, realizing $\partial P$ as a circle and letting $w$ act on this circle as a reflection with respect to some axis. A map $w$ thus extended is called a boundary involution of $f$. By definition, it reverses the orientation and satisfies $w\circ w=id_{\partial f}$, where, as always, we ignore finite sets of points.
Write $\mathrm{IETF}(\partial f)$ for the group of all Euclidean scissors automorphisms of the graph $\partial f$, not necessarily orientation preserving; these automorphisms are also known as interval exchange transformations with flips, hence the abbreviation IETF. A boundary involution $w$ introduced above is an element of $\mathrm{IETF}(\partial f)$. Let $w_{-1}$ be the boundary involution for $f^{-1}$, where the action of $w_{-1}$ is chosen to coincide with $w$ on $\partial P$. With this proviso, $w_{-1}$ is uniquely determined by $w$. Now, we can define a non-trivially iterable interval exchange transformation associated with $f^{*}$.
Definition 4.2 (A boundary quasi-commutator).
Choose boundary involutions $w$ and $w_{-1}$ as above. Define the associated boundary quasi-commutator as $(f^{*})^{-1}\circ w_{-1}\circ f^{*}\circ w$. It looks like the commutator of $f^{*}$ and $w$, except that $w_{-1}$ and $w$ are, in general, different maps defined on different graphs, hence the ‘quasi’ in the name.
The following theorem gives a useful sufficient condition for aperiodicity of $f$.
Theorem 4.3.
Let $f$ be a $\mathsf{G}$-scissors automorphism of a compact polygon $P$, where $\mathsf{G}$ is the group of all orientation preserving Euclidean isometries. If $f$ is periodic, then any associated boundary quasi-commutator has zero SAF invariant.
Proof.
Suppose that $f^{n}=id$ in the scissors automorphism group of $P$. This means that $f^{n}$ is the identity on every component of $\mathrm{dom}(f^{n})$. Cutting the graph $\partial f$ at finitely many points, one can embed it into $\partial f^{n}$. Note, however, that, in general, the latter graph can be much larger. Let $w_{n}$ be the boundary involution for $f^{n}$. In contrast to $w$, this is an element of the group $\mathrm{IETF}(\partial f^{n})$. Under the embedding of $\partial f$ into $\partial f^{n}$, both maps $w$ and $w_{-1}$ identify with certain restrictions of $w_{n}$. Hence the boundary quasi-commutator can be realized as a restriction of the actual commutator
$\sigma=[f^{*},w_{n}]:=(f^{*})^{-1}\circ w_{n}\circ f^{*}\circ w_{n}$
to an invariant subgraph $\partial f$ of $\partial f^{n}$.
Observe also that, on the complement $\partial f^{n}\setminus\partial f$, the map $\sigma$ acts as the identity, except possibly on a finite set of points, where it is undefined. Indeed, take an edge $E$ of $\partial f^{n}$ that does not belong to $\partial f$. Project it to $P$ and let $\tilde{E}$ be the image under this projection. The map $f$ is continuous at all points of $\tilde{E}$ except possibly the endpoints, which means that $w_{n}\circ f^{*}\circ w_{n}=f^{*}$ on $\tilde{E}$. Hence $\sigma=id$ almost everywhere outside $\partial f$, as claimed. It follows that the SAF invariant of $\sigma$ coincides with that of the boundary quasi-commutator.
Now it remains to show that $\mathrm{Inv}(\sigma)=0$. Knowing that $\sigma$ is a commutator of scissors congruences of $\partial f^{n}$ is not enough, since one of these congruences, $w_{n}$, reverses the orientation and therefore is not an interval exchange transformation. On the other hand,
$\mathrm{Inv}(\sigma)=-\mathrm{Inv}(f^{*})+\mathrm{Inv}(w_{n}\circ f^{*}\circ w_{n}),$
by the additivity of the SAF invariant, see Theorem 2.18. Since $f$ is periodic, so is $f^{*}$, and one has $\mathrm{Inv}(f^{*})=0$. Also, $w_{n}\circ f^{*}\circ w_{n}$ is conjugate to $f^{*}$, hence is periodic, hence has zero SAF invariant. ∎
4.2. Induced outer billiard map
Recall that $R$ is a generator of the group of all rotations of $\Pi$. It commutes with $f_{\Pi}$. For this reason, $f_{\Pi}$ induces a well defined dynamical system on the quotient space of $\mathbb{C}\setminus\Pi$ by the action of $R$. Note that $V_{0}$, a downward looking sector with vertex at $1=\zeta^{0}$ formed by continuations of two adjacent edges of $\Pi$, is a fundamental domain of $R$, and the quotient dynamics of $f_{\Pi}$ can be interpreted as a dynamics on $V_{0}$.
Definition 4.4 (Induced outer billiard map).
Let $\pi:\mathrm{dom}(f_{\Pi})\to V_{0}$ be the map sending $V_{k}:=R^{k}V_{0}$ to $V_{0}$ by $R^{-k}$, for every $k=0$, $\dots$, $N-1$. Then the induced outer billiard map $f_{\mathrm{ind}}:V_{0}\cap\mathrm{dom}(f_{\Pi}^{2})\to V_{0}$ is defined as $\pi\circ f_{\Pi}$.
As is clear from the definition, $f_{\mathrm{ind}}$ has aperiodic points if and only if $f_{\Pi}$ does. Hence it is enough to study $f_{\mathrm{ind}}$ as a piecewise isometry of the sector $V_{0}$ as well as its restrictions to bounded invariant polygons.
Remark 4.5 (Generalizations of induced outer billiards).
The presence of a regular polygon $\Pi$ is not necessary for defining the associated induced map $f_{\mathrm{ind}}$. In fact, as can be seen from the above, any choice of $\zeta$ with $|\zeta|=1$ and $\mathrm{Im}(\zeta)>0$ yields the corresponding map $f_{\mathrm{ind}}$ (depending on $\zeta$). Namely, one sets $f_{\mathrm{ind}}(z)=R^{-k_{z}}S_{0}z$, where $k_{z}$ is the smallest non-negative integer with the property $R^{-k_{z}}S_{0}w\in V_{0}$. Within the thus constructed continuous one-parameter family of dynamical systems, only discrete subset of $\zeta$ (namely, primitive roots of unity of the form $\exp(2\pi\mathbf{i}/N)$) corresponds to outer billiard maps associated with regular polygons. Note, however, that an analog of Theorem 1.1 is trivial unless $\zeta$ is a root of unity. The thus obtained one-parameter family of piecewise rotations has many features in common with the family of torus maps studied previously by Lowenstein and Vivaldi [LV16a, LV16b].
We start with describing level one pieces of $f_{\mathrm{ind}}$. The number $N$ of edges of $\Pi$ is not assumed to be even. Set $m$ to be $[\frac{N+1}{2}]-1$, where $[\frac{N+1}{2}]$ is the integer floor of $\frac{N+1}{2}$, which is consistent with our earlier use of $m$ in the even case. Let $v_{0}$ stand for the reflection (=axial symmetry) of $V_{0}$ through the vertices axis passing through point 1; the axis of $v_{0}$ is shown as a dashed vertical line in Fig. 4.
Figure 4. The domain of $f_{\mathrm{ind}}$. Left: $N=10$, middle: $N=14$, right: $N=18$ (the corresponding values of $m$ are $4$, $6$, and $8$, respectively). Dashed lines are the axes of the reflections $v_{0}$ and $v_{k}$ with $k=1$, $\dots$, $m+1$.
Lemma 4.6.
Pieces of $f_{\mathrm{ind}}$ are $Q_{k}:=v_{0}(V_{0}\cap S_{-k}V_{-k})$. Here, $k$ ranges from 1 to $m+1$. Moreover, $f_{\mathrm{ind}}$ acts as $R^{-k}\circ S_{0}$ on $Q_{k}$.
Recall that $S_{-k}$ is the reflection in the vertex of $V_{-k}$, that is, the half-turn about this vertex. Pieces $Q_{k}$ are illustrated in Fig. 4.
Proof.
First, we make several observations. The intersection $V_{0}\cap S_{-k}V_{-k}$ is nonempty if and only if $k<\frac{N+1}{2}$, that is, if $k\leqslant m+1$. It follows that $V_{0}\cap S_{-k}V_{-k}$, for $k=1$, $\dots$, $m+1$, tile the sector $V_{0}$. Another observation is that, on $V_{0}$, the map $f_{\Pi}=S_{0}$ acts as $u_{0}\circ v_{0}$, where $u_{0}$ is the reflection in the real line, that is, the complex conjugation. Observe, finally, that $u_{0}$ takes the sector $S_{-k}V_{-k}$ to the sector $V_{k}$. Indeed, since $V_{-k}$ has vertex $\zeta^{-k}$, the sector $u_{0}S_{-k}V_{-k}$ has vertex $\zeta^{k}=u_{0}(\zeta^{-k})$. Now, it only remains to figure out which of the two sectors with vertex $\zeta^{k}$ bounded by support lines of $\Pi$ is represented by $u_{0}V_{-k}$ and which by $u_{0}S_{-k}V_{-k}$; this is straighforward by noticing that $V_{-k}$ and $u_{0}S_{-k}V_{-k}$ are oriented clockwise with respect to the origin.
We proceed with the proof of the lemma: $S_{0}=u_{0}\circ v_{0}$ takes the polygon $Q_{k}$ to $u_{0}(V_{0}\cap S_{-k}V_{-k})\subset u_{0}S_{-k}V_{-k}=V_{k}.$ Hence, $\pi$ acts as $R^{-k}$ on $S_{0}(Q_{k})$, and $f_{\mathrm{ind}}=R^{-k}\circ S_{0}$ on $Q_{k}$. ∎
Next, it will be convenient to represent $f_{\mathrm{ind}}$ as a composition of two involutions, one of which is $v_{0}$.
Lemma 4.7.
The map $v_{0}\circ f_{\mathrm{ind}}$ acts on each of the pieces $Q_{k}$ as a reflection $v_{k}$ in a symmetry axis of $Q_{k}$. Hence, $f_{\mathrm{ind}}=v_{0}\circ v_{k}$ on $Q_{k}$.
It follows, in particular, that each $Q_{k}$ has an axis of symmetry; these axes are shown as dashed lines in Fig. 4.
Proof.
As is easily seen, there is an axial symmetry of $\Pi$ that interchanges the vertices $1$ and $\zeta^{-k}$ of $\Pi$. The same reflection $\tilde{v}_{k}$ interchanges the sectors $V_{0}$ and $S_{-k}V_{-k}$. Hence, it induces an axial symmetry of the intersection $V_{0}\cap S_{-k}V_{-k}$ and, therefore, also of $Q_{k}=v_{0}(V_{0}\cap S_{-k}V_{-k})$ (see Lemma 4.6). Let $v_{k}$ be the reflection in the latter symmetry axis of $Q_{k}$, that is, $v_{k}:=v_{0}\circ\tilde{v}_{k}\circ v_{0}$. It now suffices to show that $f_{\mathrm{ind}}$ acts as $\tilde{v}_{k}\circ v_{0}$ on $Q_{k}$.
By Lemma 4.6, the map $f_{\mathrm{ind}}$ acts as $R^{-k}\circ S_{0}$ on $Q_{k}$. Extend $\tilde{v}_{k}$ and $v_{0}$ to global Euclidean reflections, while keeping the same notation. The map $v_{0}$ takes $V_{0}$ to itself, while $\tilde{v}_{k}$ sends it to $S_{-k}V_{-k}$. On the other hand, $S_{0}$ takes $V_{0}$ to $S_{0}V_{0}$, the sector symmetric to $V_{0}$ with respect to its vertex $1$, and $R^{-k}$ sends $S_{0}V_{0}$ to $S_{-k}V_{-k}$. Since both $f_{\mathrm{ind}}$ and $\tilde{v}_{k}\circ v_{0}$ take the sector $V_{0}$ to the same sector and since both maps preserve the orientation, these two maps must coincide. ∎
4.3. Geometry of the pieces $Q_{k}$
In Section 4.3, we assume that $N=2m+2$ is twice odd, and that $m$ is even, that is, $N\equiv 2$ modulo 4. Below, we compute geometric parameters of the pieces $Q_{k}$ such as the angles and the sidelengths, i.e., lengths of the edges. Observe that the pieces $Q_{1}$ through $Q_{m/2}$ form a triangle $U$ that is invariant under $f_{\mathrm{ind}}$ in the sense that, for every piece $Q_{k}\subset U$, we also have $f_{\mathrm{ind}}(Q_{k})\subset U$. The triangle $U$ is isosceles with horizontal base; the existence of such a triangle depends on our assumption that $m$ is even. In particular, the restriction $f_{U}$ of $f_{\mathrm{ind}}$ to $U$ is an orientation preserving scissors automorphism of $U$.
Denote the vertices of the level 1 pieces of $f_{U}$ that lie on the left edge of $U$ by $a_{1}$, $\dots$, $a_{m/2}$; the numeration is downward. Similarly, let $b_{1}$, $\dots$, $b_{m/2}$ be the vertices on the right edge of $U$; see Fig. 5. In particular, $Q_{1}$ is the triangle with vertices $\zeta^{0}$, $a_{1}$, $b_{1}$, and each $Q_{k}$ with $k>1$ is a quadrilateral with vertices $a_{k}$, $b_{k}$, $b_{k-1}$, $a_{k-1}$, where the vertices are always listed in a counterclockwise order. Note also that the reflection $v_{k}$ acts on $Q_{k}$ by fixing the vertices $a_{k}$, $b_{k-1}$ while interchanging $a_{k-1}$, $b_{k}$.
Figure 5. Pieces $Q_{k}$ of $U$. Symmetry axes of $U$ and of all $Q_{k}$ are shown as dashed lines. This figure is simply a zoom-in of Fig. 4.
Write $\lambda_{k}$ for the length of $[a_{k};b_{k}]$, where $k$ ranges from 1 to $m/2$. By the symmetry of the pieces $Q_{k}$, we also have
$|a_{k-1}-a_{k}|=\lambda_{k}=|b_{k}-b_{k+1}|,$
where the left equality holds for $k>1$ as stated and for $k=1$ after $a_{0}$ is replaced with 1; the right equality holds for $k<m/2$. Recall from Section 3.1 that $\lambda_{0}=2\sin\frac{\theta}{2}$ is the length of each edge of $\Pi$, which is equal to $|\zeta^{0}-b_{1}|$. Elementary geometry considerations allow to compute all angles of the pieces $Q_{k}$, and then the lengths $\lambda_{k}$ for all $k=1$, $\dots$, $m/2$.
Lemma 4.8.
The triangle $Q_{1}$ has the following angles and sidelengths
$\angle b_{1}=\theta,\ \angle a_{1}=\pi-2\theta,\ \angle\zeta^{0}=\theta,\ |\zeta^{0}b_{1}|=\lambda_{0},\ |a_{1}b_{1}|=|a_{1}\zeta^{0}|=\lambda_{1}.$
For $k=2$, $\dots$, $m/2$, the quadrilateral $Q_{k}$ has angles
$\angle b_{k-1}=\pi-(k-1)\theta,\ \angle a_{k}=\pi-(k+1)\theta,\ \angle b_{k}=\angle a_{k-1}=k\theta$
and sidelengths $|b_{k-1}b_{k}|=|b_{k-1}a_{k-1}|=\lambda_{k-1}$, $|b_{k}a_{k}|=|a_{k-1}a_{k}|=\lambda_{k}$. Here, $\lambda_{0}=2\sin\frac{\theta}{2}$ and, for $k=1$, $\dots$, $m-1$,
$\lambda_{k}=\frac{\sin\theta}{2\cos\frac{k\theta}{2}\cos\frac{(k+1)\theta}{2}}=\frac{\lambda_{0}\zeta^{k}(1+\zeta)}{(1+\zeta^{k})(1+\zeta^{k+1})},\quad\zeta:=e^{2\pi\mathbf{i}/N}.$
Proof.
The claim about the angles of $Q_{1}$ is straighforward. Next, consider the quadrilateral $Q_{2}$. Since the angle $\angle a_{1}$ of $Q_{2}$ is an external angle of $Q_{1}$, it can be computed as the sum of the two non-adjacent internal angles, that is, $\angle a_{1}=\theta+\theta=2\theta$. By the same reason, the angle $\angle b_{1}$ of $Q_{2}$ equals $\theta+(\pi-\theta)=\pi-2\theta$. Also, $\angle a_{1}=\angle b_{2}$ by symmetry. Finally, the fourth angle $\angle a_{2}$ of the quadrilateral $Q_{2}$ can be recovered from the three angles already computed. Now, assume by induction that the claim about the angles of $Q_{k-1}$ holds for some $k>2$, and look at the angles of $Q_{k}$. The argument is similar to the above. Firstly, the angle $\angle a_{k-1}$ of $Q_{k}$ is an external angle of the triangle $a_{k-1}b_{k-1}\zeta^{0}$. As such, it equals the sum of the two non-adjacent internal angles of the same triangle, namely, to $(k-1)\theta+\theta=k\theta$. Here, we used that the angle $\angle a_{k-1}b_{k-1}\zeta^{0}$ is $(k-1)\theta$ by the induction hypothesis. Secondly, the angle $\angle b_{k-1}$ can also be computed as the sum of the two non-adjacent angles of the triangle $a_{k-1}b_{k-1}\zeta^{0}$, that is, as $\theta+(\pi-k\theta)=\pi-(k-1)\theta$. It remains to note that $\angle b_{k}=\angle a_{k-1}$ in $Q_{k}$ by symmetry, and that the fourth angle of $Q_{k}$ can be recovered by the remaining three, already computed. We have thus proved the claims about the angles of all $Q_{k}$.
Let us compute the sidelengths. First, note that $\lambda_{0}=2\lambda_{1}\cos\theta$, as follows from studying the isosceles triangle $Q_{1}$. Hence, $\lambda_{1}=\frac{\sin\frac{\theta}{2}}{\cos\theta}$. A recurrence relation for $\lambda_{k}$ can be obtained from the Law of Sines. Namely, by the Law of Sines applied to the triangle $b_{k}a_{k+1}a_{k}$, we obtain
$\frac{\lambda_{k}}{\sin\frac{\pi-(k+2)\theta}{2}}=\frac{\lambda_{k+1}}{\sin\frac{\pi-k\theta}{2}}\quad\Longrightarrow\quad\lambda_{k+1}=\frac{\cos\frac{k\theta}{2}}{\cos\frac{(k+2)\theta}{2}}\lambda_{k}.$
Solving this recurrence relation, together with the initial condition $\lambda_{1}=\frac{\sin\frac{\theta}{2}}{\cos\theta}$, we obtain that
$\lambda_{2}=\frac{\sin\theta}{2\cos\theta\,\cos\frac{3\theta}{2}},\quad\lambda_{k}=\frac{\sin\theta}{2\cos\frac{k\theta}{2}\cos\frac{(k+1)\theta}{2}}.$
We can also express $\lambda_{k}$ via $\lambda_{0}$ and $\zeta$ as follows:
$\lambda_{k}=\frac{\lambda_{0}(\zeta^{\frac{1}{2}}+\zeta^{-\frac{1}{2}})}{(\zeta^{\frac{k+1}{2}}+\zeta^{-\frac{k+1}{2}})(\zeta^{\frac{k}{2}}+\zeta^{\frac{-k}{2}})}=\frac{\lambda_{0}\zeta^{k}(1+\zeta)}{(1+\zeta^{k})(1+\zeta^{k+1})}.$
Here, by $\zeta^{t}$ we mean $\exp(2\pi\mathbf{i}t\theta)$, even if $t$ is not integer. ∎
4.4. Computing boundary SAF invariants
A boundary SAF invariant of an orientation preserving scissors automorphism $f$ is defined as the SAF invariant of a boundary quasi-commutator of $f$. Computing boundary SAF invariants for $f$ helps establishing aperiodicity. Below, we give a specific example of such a computation. Namely, consider the induced billiard map $f_{\mathrm{ind}}$ in the case when $N=2m+2$ is even but not divisible by 4; in this case, $m$ is also even. As follows from the above, $U$ is an invariant triangle of $f_{\mathrm{ind}}$, and $f_{U}$ stands for the restriction of $f_{\mathrm{ind}}$ to $U$. We will compute a boundary SAF invariant of $f_{U}$. First, one needs to define the boundary involution on the external boundary of $U$. A standard choice is the axial symmety of $U$, that is, the reflection $v_{0}$ of $U$ in the vertical axis passing through point $1$. Both the boundary involution and the boundary quasi-commutator are now well defined for $f_{U}$.
The boundary quasi-commutator $\sigma$ of $f_{U}$ is what we need to compute. Write $u$ for the reflection $v_{0}$ restricted to $U$, and let $v$ be the piecewise isometry of $U$ acting as $v_{k}$ on each piece $Q_{k}$, for $k=1$, $\dots$, $m/2$. By Lemma 4.7, the map $f_{U}$ can be represented as $u\circ v$. From this and from the observation that both $u$ and $v$ act on their respective domains as involutions, it follows that the inverse map $f^{-1}_{U}$ is conjugate to $f_{U}$ by means of $u$:
$f^{-1}_{U}=u\circ f_{U}\circ u,$
a property that plays a definite role in what follows. Let $u^{*}$ and $v^{*}$ be the induced by $u$ and $v$, respectively, maps of $\partial f$; the former sending $\partial f$ to $\partial f^{-1}$, and the latter — to $\partial f$ itself. Also, write $f^{*}$ instead of $f^{*}_{\mathrm{ind}}$ to lighten the notation. Since $f_{\mathrm{ind}}=u\circ v$ by Lemma 4.7, one also has $f^{*}=u^{*}\circ v^{*}$.
Lemma 4.9.
The boundary quasi-commutator $\sigma$ of $f^{*}$ is the second iterate of $v^{*}\circ w$. In particular, $\mathrm{Inv}(\sigma)=2\,\mathrm{Inv}(v^{*}\circ w)$.
Proof.
A key observation is that $u^{*}\circ w_{-1}\circ u^{*}=w$, which follows from the fact that $f^{-1}=u\circ f\circ u$. Using this observation, we can write
$\sigma=(f^{*})^{-1}\circ w_{-1}\circ f^{*}\circ w=(v^{*}\circ u^{*})\circ w_{-1}\circ(u^{*}\circ v^{*})\circ w=(v^{*}\circ w)^{2}.$
Here, the first equality uses the representation of $f^{*}$ as $v^{*}\circ u^{*}$, and the second equality uses the observation just made. ∎
It remains to compute the SAF invariant of the map $v^{*}\circ w$. Recall that the outer boundary $\partial_{out}f_{U}$ of of $f_{U}$ is the part of $\partial f_{U}$ that projects onto the boundary of $U$. On $\partial_{out}f_{U}$, this projection is one-to-one except on finitely many points, namely, except on vertices of the subgraph $\partial_{out}f$ of the graph $\partial f$. Denote by $\mathcal{R}_{v^{*}\circ w}$ the first return map of $\partial_{out}f_{U}$ with respect to $v^{*}\circ w$. For completeness, we recall the general concept of a first return map below.
Definition 4.10 (First return map).
Let $F$ be a partially defined map of a set $D$ to itself, that is, $F:\mathrm{dom}(F)\to D$, where $\mathrm{dom}(F)\subset D$. For any subset $A\subset D$ and any point $x\in A$, the first return time $n_{x}$ to $A$ is defined as the smallest positive integer $n$ with the property $F^{n}(x)\in A$ or infinity if such $n>0$ does not exist. By definition, the domain of the first return map $\mathcal{R}_{A}F$ consists of all points $x\in A$ such that $n_{x}<\infty$. Every $x\in\mathrm{dom}(\mathcal{R}_{A}F)$ is mapped to the point $\mathcal{R}_{A}F(x)=F^{n_{x}}(x)$.
In our case, $D=\partial f_{U}$, $A=\partial_{out}f_{U}$, and $F=v^{*}\circ w$. The first return map $\mathcal{R}_{out}:=\mathcal{R}_{A}F$ can therefore be viewed as an interval exchange transformation of $\partial U$, upon the identification between $\partial U$ and $\partial_{out}f_{U}$. We can explicitly describe the map $\mathcal{R}_{out}$.
Consider a homeomorphism between $\partial U$ and a circle that preserves Euclidean lengths. If a self-map $h$ of $\partial U$ is conjugate to a rotation by angle $\alpha$ under such a homeomorphism, then we say that $h$ is a rotation of $\partial U$ by angle $\alpha$. Here, angles are viewed as elements of $\mathbb{R}/\mathbb{Z}$, in particular, the angular measure of the full circle is 1 rather than $2\pi$. Similarly, if $h$ is conjugate to a reflection of the circle in some diameter chord of it, then we say $h$ is a reflection of $\partial U$. Evidently, every reflection of $\partial U$ has a pair of distinct fixed points.
Lemma 4.11.
The map $\mathcal{R}_{out}$ is the rotation of $\partial U$ by angle
$\alpha=\frac{\sin\frac{\theta}{2}}{1+\sin\frac{\theta}{2}}\in\mathbb{R}/\mathbb{Z}.$
Proof.
Recall that $\overline{v}_{i}$ is the continuous extension of $v_{i}$ to $\overline{Q}_{i}$, the closure of $Q_{i}$. Similarly, write $\overline{u}$ for the continuous extension of $u$ to $\overline{U}$.
Define a map $\rho\in\mathrm{IETF}(\partial U)$ as follows. The domain of $\rho$ will be the complement in $\partial U$ of finitely many points; namely, all vertices of $U$ and all vertices of all $Q_{i}$, the components of $\mathrm{dom}(f_{U})$. Now, take a point $z\in\mathrm{dom}(\rho)$; then $z$ lies on the boundary of a (necessarily unique) component $Q_{i}$. There are two cases: either $\overline{v}_{i}(z)$ lies in $\partial U$, or not (the latter case means that $\overline{v}_{i}(z)$ is a common boundary point of two pieces of $f_{U}$). In the former case, set $\rho(z):=\overline{v}_{i}(z)$. Consider the latter case: there is a unique $Q_{j}\neq Q_{i}$ such that $\overline{v}_{i}(z)\in\partial Q_{j}$. Set $\rho(z)=\overline{v}_{j}\circ\overline{v}_{i}(z)$; this point necessarily lies in $\partial U$. In terms of the dynamics on the graph $\partial f_{U}$, the map corresponding to $\rho$ can be written as $v^{*}\circ w\circ v^{*}$ in the latter case and as $v^{*}$ in the former case.
It is not hard to describe the map $\rho$ in our specific situation: $\rho$ acts on $\partial U$ as a reflection with fixed point $a_{m/2}$; the other fixed point is the midpoint $b_{0}$ of the segment $[1;b_{1}]$, that is, the point where the symmetry axis of $Q_{1}$ crosses the edge $[1;b_{1}]$. Recall that this is an orientation reversing isometry of $\partial U$ (with respect to the length metric) which is conjugate to a Euclidean reflection under any isometric homeomorphism between $\partial U$ and a circle. Explicitly, $\rho$ interchanges the following pairs of straight line segments:
$[b_{0};1]\leftrightarrow[b_{0};b_{1}],$ $\quad[b_{1};b_{2}]\leftrightarrow[1;a_{1}],$
$[b_{2};b_{3}]\leftrightarrow[a_{1};a_{2}],$ $\quad\dots,\quad[b_{m/2};a_{m/2}]\leftrightarrow[a_{m/2-1};a_{m/2}].$
Here, the map $\rho$ also preserves the order of the endpoints. For example, $[b_{1};b_{2}]$ is mapped under $v^{*}_{2}$ to $[b_{1};a_{1}]$, and the latter segment is mapped under $v^{*}_{1}$ to $[1;a_{1}]$.
Recall that $\overline{u}$ also acts as a reflection of $\partial U$; its fixed points are $1$ and the midpoint of $[a_{m/2};b_{m/2}]$. As a composition of two reflections, $\mathcal{R}_{out}=\rho\circ u^{*}$ is necessarily a rotation by some angle; it remains only to compute the angle $\alpha$ of this rotation. Note that the point $b_{1}$ is mapped to the point $1$; hence
$\alpha=\frac{\lambda_{0}}{\mathrm{Len}(\partial U)}.$
Here, the denominator is the perimeter of $U$, and the numerator is the signed distance from $b_{1}$ to $1$ in the boundary $\partial U$ oriented counterclockwise. Note that $U$ is an isosceles triangle with base $[a_{m/2};b_{m/2}]$ and the opposite angle $\theta$. Using the notation of Lemma 4.8, the length of the base is
$\lambda_{m/2}=\lambda_{0}\frac{\zeta^{\frac{m}{2}}(1+\zeta)}{(1+\zeta^{\frac{m}{2}})(1+\zeta^{\frac{m}{2}+1})}=\lambda_{0}.$
Here, we used that $\zeta^{m/2}=i\zeta^{-1/2}$, which follows from $\zeta^{m+1}=-1$. Therefore,
$\mathrm{Len}(\partial U)=\lambda_{0}\frac{1+\sin\frac{\theta}{2}}{\sin\frac{\theta}{2}}.$
The desired formula for the angle $\alpha$ follows. ∎
It remains to find a relationship between the SAF invariants of the interval exchange maps $v^{*}\circ w$ and $\mathcal{R}_{out}$. Recall that an interval swap map was defined in Section 2.3, see Example 2.12. By definition of the SAF invariant, it is immediate that $\mathrm{Inv}(s)=0$ for every interval swap map.
Lemma 4.12.
Write $\mathcal{R}$ for the self-map of $\partial f$ acting as $\mathcal{R}_{out}$ on $\partial_{out}f$ and as the identity elsewhere. The map $\mathcal{R}\circ(v^{*}\circ w)^{-1}$ is a composition of interval swap maps, hence $\mathrm{Inv}(\mathcal{R}_{out})=\mathrm{Inv}(\mathcal{R})=\mathrm{Inv}(v^{*}\circ w)$.
An essential property of the map $v^{*}$ we will be using is the following: if $x\in\partial_{in}f$, then $v^{*}(x)\in\partial_{out}f$. In particular, the first return time to $\partial_{out}f$ of any point of $\partial_{out}f$ under $v^{*}\circ w$ is either one or two. This is, in fact, the only specific property of $f$ we use in the proof that follows; Lemma 4.12 therefore admits a significant generalization.
Proof.
For every edge $I$ of the graph $\partial_{in}f$, consider the interval swap map $s_{I}$ interchanging $I$ with $v^{*}\circ w(I)$. Evidently, $s_{I}$ acts as $v^{*}\circ w$ on $I$ and as $w\circ v^{*}$ on $v^{*}\circ w(I)$. The maps $s_{I}$ for different $I$s have disjoint supports, ignoring finite sets of points, as always. Hence we may talk about the composition of the maps $s_{I}$, not caring about the order. Denote this composition by $s$; it is a composition of interval swap maps by construction.
We claim that $\mathcal{R}=v^{*}\circ s\circ w$. Evidently, it suffices to check this claim on any point $x$ inside an edge $E$ of $\partial f$, i.e., when $x\in\partial f$ not $x$ is not a vertex of $\partial f$. Consider cases. If $x\in\partial_{out}f$ and $v^{*}\circ w(x)\in\partial_{out}f$, then, by definition of the first return map, $\mathcal{R}(x)=v^{*}\circ w(x)$, and $s(w(x))=w(x)$, which implies the desired claim in this case. Next, if $x\in\partial_{out}f$ and $v^{*}\circ w(x)\in\partial_{in}f$, then, on the one hand, $\mathcal{R}(x)=(v^{*}\circ w)^{2}(x)$, by the property mentioned above, and, on the other hand,
$v^{*}\circ s\circ w(x)=v^{*}\circ w\circ v^{*}\circ w(x)=\mathcal{R}(x),$
since $w(x)\in v^{*}\circ w(I)$ for some $I\subset\partial_{in}f$, in this case. Finally, if $x\in\partial_{in}f$, then $s$ acts as $v^{*}\circ w$ on $w(x)$, and
$v^{*}\circ s\circ w(x)=v^{*}\circ v^{*}\circ w\circ w(x)=x=\mathcal{R}(x),$
which also verifies the claim in the last case.
Thus, in all cases, the map $\mathcal{R}$ can be written as
$\mathcal{R}=v^{*}\circ s\circ w=(v^{*}\circ s\circ v^{*})\circ(v^{*}\circ w)=s^{*}\circ(v^{*}\circ w),$
where $s^{*}:=v^{*}\circ s\circ v^{*}$ is conjugate to $s$ by means of $v^{*}$, in particular, $s^{*}$ is also a composition of interval swap maps. ∎
Combining all the results proved so far in Section 4.4, we obtain an expression for the SAF invariant of $\sigma$, the boundary quasi-commutator of $f^{*}_{U}$.
Theorem 4.13.
The boundary quasi-commutator $\sigma$ of $f^{*}_{U}$ has SAF invariant
$\mathrm{Inv}(\sigma)=2\wedge_{\mathbb{Q}}\alpha\neq 0.$
The angle $\alpha\in\mathbb{R}/\mathbb{Z}$ is that introduced in Lemma 4.12.
Proof.
This is a matter of gathering the already obtained results. Firstly, by Lemma 4.9, the invariant of $\sigma$ is twice the invariant of $v^{*}\circ w$. Secondly, by Lemma 4.12, the latter invariant coincides with $\mathrm{Inv}(\mathcal{R}_{out})$. Lastly, $\mathcal{R}_{out}$ is the rotation of $\partial U$ by angle $\alpha$. By definition, it is conjugate to the rotation by angle $\alpha$ of the unit circle $\mathbb{R}/\mathbb{Z}$ or, which is the same, the map $x\mapsto\{x+\alpha\}$ on $[0;1)$. It is easy to compute the SAF invariant of the latter map:
$(1-\alpha)\otimes_{\mathbb{Q}}\alpha+\alpha\otimes_{\mathbb{Q}}(\alpha-1)=1\wedge_{\mathbb{Q}}\alpha.$
The desired formula follows. ∎
Theorem 4.13 implies again, by a different method, a part of Theorem 3.8, namely, that the outer billiard map on a regular $(2m+2)$-gon for $m$ even has non-periodic restriction to some bounded invariant polygon.
5. Topological dynamics
Section 5 considers a very general setup of topological dynamics, namely, an arbitrary compact metric space $X$ and an arbitrary homeomorphism $h:X\to X$. Iterates of $h$ generate a dynamical system on $X$ whose properties are studied below. A specific metric on $X$ is never used; what we need is only that $X$ is metrizable. Notions of topological dynamics are developed in a similar, or even greater, generality in [Aki93]. We do not claim (much) originality in Section 5, however, we were unable to find convenient references for some of the results that follow.
5.1. Boundedness
Write $\mathbb{R}_{>c}$ for the set of reals $t$ such that $t>c$. Consider a lower semicontinuous function $\xi:X\to\mathbb{R}_{>0}$. Recall that $\xi$ being lower semicontinuous means that $\xi^{-1}(\mathbb{R}_{>c})$ is open for every $c\in\mathbb{R}_{>0}$.
Definition 5.1 (Quasi-invariance).
Say that $\xi$ is quasi-invariant under $h$ if there is an increasing homeomorphism $\Xi:\mathbb{R}_{>0}\to\mathbb{R}_{>0}$ with the property that $\xi\circ h<\Xi\circ\xi$ and $\xi\circ h^{-1}<\Xi\circ\xi$. The homeomorphism $\Xi$ is called the modulus of quasi-invariance.
An example of a quasi-invariant function is an invariant function, i.e., a function $\xi$ satisfying the identity $\xi\circ h=\xi$.
Theorem 5.2.
Let $X$ be a nonempty compact metric space, and $h:X\to X$ be a homeomorphism. Then any quasi-invariant under $h$ lower semicontinuous function $\xi:X\to\mathbb{R}_{>0}$ is bounded on some nonempty compact $h$-invariant subset of $X$.
Assume, by way of contradiction, that $\xi$ is unbounded on every compact $h$-invariant subset of $X$. We keep this assumption as our standing assumption throughout the proof of Theorem 5.2. Say that $\xi$ is locally unbounded at a point $x\in X$ if it is unbounded on every neighborhood of $x$. The proof of Theorem 5.2 will use a transfinite sequence of compact invariant subsets of $X$ defined below. Define $Y_{0}$ as $X$. For any ordinal number $\alpha>0$, finite or infinite, assume by transfinite induction that $Y_{\alpha}$ is already defined, and define $Y_{\alpha+1}$ as the set of all points of $Y_{\alpha}$, at which the restriction $\xi_{\alpha}:=\xi|_{Y_{\alpha}}$ is locally unbounded. It remains to define $Y_{\alpha}$ for a limit ordinal $\alpha$ assuming that $Y_{\beta}$ is already defined for all $\beta<\alpha$. Let $Y_{\alpha}$ simply be the intersection of all such $Y_{\beta}$. We have thus defined a transfinite sequence of subsets $Y_{\alpha}\subset X$.
Lemma 5.3.
Every subset $Y_{\alpha}$ is compact, non-empty, and $h$-invariant.
Recall that $Y_{\alpha}$ being $h$-invariant means $h(Y_{\alpha})=Y_{\alpha}$.
Proof.
For any compact metric space $Y$ and any function $\eta:Y\to\mathbb{R}_{>0}$, the set of all points of $Y$ where $\eta$ is locally unbounded is compact. Also, the intersection of any family of compact sets is compact. It follows that all $Y_{\alpha}$ are compact by a straightforward transfinite induction.
Let us now show that $Y_{\alpha+1}$ is nonempty assuming that $Y_{\alpha}\neq\varnothing$ and that $Y_{\alpha}$ is $h$-invariant. Note that $\xi_{\alpha}$ is unbounded on $Y_{\alpha}$ by our standing assumption that there is no invariant compact subset on which $\xi$ is bounded. By compactness, it follows that $\xi_{\alpha}$ is locally unbounded at some points of $Y_{\alpha}$, that is, $Y_{\alpha+1}$ is nonempty. Also, the intersection of a nested transfinite sequence of nonempty compact sets is nonempty. It follows that all $Y_{\alpha}$ are nonempty by transfinite induction.
Since $\xi$ is quasi-invariant under $h$, it follows that $\xi$ being locally unbounded at $x\in X$ implies it also being locally unbounded at $h(x)$ and $h^{-1}(x)$. From here and the obvious observation that the intersection of $h$-invariant sets is also $h$-invariant, the $h$-invariance of all $Y_{\alpha}$ easily follows. ∎
The proof of Theorem 5.2 now becomes a straightforward consequence of Baire’s theorem (see, e.g. Section 7.3 of [KF75]): if $Y$ is a complete metric space, then any countable union of closed nowhere dense subsets of $Y$ has empty interior, in particular, cannot be $Y$.
Proof of Theorem 5.2.
Note that the transfinite sequence $Y_{\alpha}$ stabilizes: $Y_{\alpha}=Y_{\beta}$ for all sufficiently large $\alpha$ and $\beta$, for example, for $\alpha$ and $\beta$ having the cardinal number bigger than than of $X$. Set $Y$ to be $Y_{\alpha}$ for such a large ordinal $\alpha$. Then $\xi_{Y}:=\xi_{\alpha}$ is locally unbounded at every point of $Y$. Also, by Lemma 5.3, the set $Y$ is compact and nonempty. Consider the sets $Y^{\leqslant n}:=\{y\in Y\mid\xi(y)\leqslant n\}$ for all positive integer values of $n$. By lower semicontinuity of $\xi$, each set $Y^{\leqslant n}$ is closed, and, since $\xi$ is nowere locally bounded on $Y$, it is nowhere dense. As follows from Baire’s theorem, the union of $Y^{\leqslant n}$ is nowhere dense; but the union is $Y$ itself, a contradiction. ∎
Theorem 5.2 motivates the following definition.
Definition 5.4 (The $\xi$-accumulation set).
Let $X$, $h$, and $\xi$ be as in Theorem 5.2. Instead of $Y_{\alpha}$, we will now write $X^{\xi}_{\alpha}$. By definition, $X^{\xi}_{0}=X$, and $X^{\xi}_{\alpha+1}$ is defined as the subset of $X^{\xi}_{\alpha}$ consisting of all points at which the restriction of $\xi$ to $X^{\xi}_{\alpha}$ is locally unbounded. Finally, for any limit ordinal $\alpha$, the subset $X^{\xi}_{\alpha}$ is defined as the intersection of all $X^{\xi}_{\beta}$ with $\beta<\alpha$. All $X^{\xi}_{\alpha}$ are compact subsets of $X$. From Theorem 5.2, it follows that $X^{\xi}_{\alpha+1}=\varnothing$ but $X_{\alpha}\neq\varnothing$ for some $\alpha$; hence $\xi$ is bounded on $X^{\xi}_{\alpha}$. This subset $X^{\xi}_{\alpha}$ is called the $\xi$-accumulation set of $X$ and is denoted by $X^{\xi}$. Note that $\xi$ is bounded on $X^{\xi}$ but is unbounded in a neighborhood (in $X$) of any point of $X^{\xi}$.
The $\xi$-accumulation set $X$ depends monotonely on $X$ with respect to inclusion, as the following lemma shows.
Lemma 5.5.
Suppose that $Y\subset X$ is a compact $h$-invariant subset of $X$. Then $Y^{\xi}\subset X^{\xi}$, where $Y^{\xi}$ denotes the $\xi|_{Y}$-accumulation set of $Y$.
Proof.
It suffices to show that $Y^{\xi}_{\alpha}\subset X^{\xi}_{\alpha}$ by transfinite induction on $\alpha$. The induction base is given as an assumption. Next, if $Y^{\xi}_{\alpha}\subset X^{\xi}_{\alpha}$, then any point of $Y^{\xi}_{\alpha}$, where the restriction of $\xi$ to $Y^{\xi}_{\alpha}$ is locally unbounded, is also a point of $X^{\xi}_{\alpha}$, at which the restriction of $\xi$ to $X^{\xi}_{\alpha}$ is locally unbounded. Hence, $Y^{\xi}_{\alpha+1}\subset X^{\xi}_{\alpha+1}$. Passing to the intersections also preserves the inclusion, and the lemma now follows by transfinite induction. ∎
5.2. Centers
If, in Theorem 5.2, the homeomorphism $h$ has periodic points, then the conclusion of the theorem becomes trivial. Thus, the most interesting case is when $h:X\to X$ has no periodic points; this is our standing assumption for Section 5.2. Below, we restate some known results from topological dynamics, cf. [KH95, Section 3.3].
Theorem 5.6.
Suppose that $X$ is a nonempty compact metric space, and $h:X\to X$ is a homeomorphism without periodic points. Then $X$ contains a nonempty perfect invariant subset $Y$. Moreover, points of $Y$ are nonwandering under the iterates of $h|_{Y}$ in the following sense: for every $y\in Y$ and every neighborhood $O$ of $y$ in $Y$, there is $n>0$ with $f^{n}(O)\cap O\neq\varnothing$.
Note that $O$ is a neighborhood of $y$ in $Y$ rather than in $X$. A point $y\in Y$ being nonwandering means that nearby points return to arbitrarily small neighborhoods of $y$ under the iterates of $h$. Recall that a perfect set is by definition a compact set without isolated points.
To prove Theorem 5.6, we define a transfinite sequence of $h$-invariant compact subsets $X_{\alpha}$ labeled by ordinal numbers $\alpha$ as follows. Set $X_{0}=X$. Next, if $\alpha$ is any ordinal, and assuming by induction that $X_{\alpha}$ is already defined, define $X_{\alpha+1}$ as the set of all nonwandering points of $h|_{X_{\alpha}}$. By definition, $X_{\alpha+1}\subset X_{\alpha}$. It remains to define $X_{\alpha}$ for a limit ordinal $\alpha$ provided that $X_{\beta}$ are already defined for all $\beta<\alpha$. We set $X_{\alpha}$ to be the intersection of $X_{\beta}$ for all $\beta<\alpha$, in this case. Thus, all sets $X_{\alpha}$ are defined by transfinite induction. As is easy to see, the transfinite sequence $X_{\alpha}$ eventually stabilizes: $X_{\alpha}=X_{\beta}$ for $\alpha$ sufficiently large (say, such that the cardinality of $\alpha$ is larger than that of $X$) and all $\beta>\alpha$.
Lemma 5.7.
Any set $X_{\alpha}$ is closed, hence compact, and nonempty.
Proof.
The set $X_{0}$ is closed by definition, and the intersection of closed sets is closed. It is therefore enough to show that $X_{\alpha+1}$ is closed assuming that so is $X_{\alpha}$. Consider a sequence $x_{n}\in X_{\alpha+1}$ converging to some point $x\in X$. We want to prove that $x\in X_{\alpha+1}$. Pick a neighborhood $O_{x}$ of $x$ in $X_{\alpha}$; then all $x_{n}$ with sufficiently large $n$ lie in $O_{x}$. By definition of $X_{\alpha+1}$ and since $O_{x}$ is also a neighborhood of $x_{n}\in X_{\alpha+1}$ with large $n$, there is a point $y\in O_{x}\cap X_{\alpha}$ and a positive integer $m$ such that $h^{m}(y)\in O_{x}$. The existence of such $y$ proves that $x\in X_{\alpha+1}$, as claimed. We obtain that all $X_{\alpha}$ are closed, hence compact.
Let us now prove that they are nonempty; we will use transfinite induction again. Induction base $\alpha=0$ is clear: by our assumption, $X_{0}=X$ is nonempty. Assume that $X_{\alpha}\neq\varnothing$ and look for a point in $X_{\alpha+1}$. Choose any $x\in X_{\alpha}$ and let $y$ be any accumulation point of the forward orbit $(h^{n}(x))$, that is, a limit of some converging subsequence $(h^{n_{k}}(x))_{k}$. Let $O_{y}$ be any neighborhood of $y$; then $h^{n_{k}}(x)\in O_{y}$ for all sufficiently large $k$. Note, in particular, that $z:=h^{n_{k}}(x)\in O_{y}\cap X_{\alpha}$ and also
$h^{n_{k+1}}(x)=h^{n_{k+1}-{n_{k}}}(z)\in O_{y},$
which proves that $y\in X_{\alpha+1}$. Finally, if $X_{\beta}\neq\varnothing$ for all $\beta<\alpha$, then the intersection $X_{\alpha}$ of all such $X_{\beta}$ is nonempty as the intersection of a nested transfinite sequence of nonempty compact sets. ∎
The transfinite sequence $(X_{\alpha})$ allows to prove Theorem 5.6.
Proof of Theorem 5.6.
If ordinals $\alpha$ are sufficiently large, then all $X_{\alpha}$ are the same; denote them by $Y$. By Lemma 5.7, the set $Y$ is compact and nonempty. Also, by definition of $X_{\alpha}$ and since $X_{\alpha+1}=X_{\alpha}$, the set $Y$ satisfies the following property: for every $x\in Y$ and any neighborhood $O_{x}$ of $x$, there is a point $y\in O_{x}$ and a positive integer $n$ with $h^{n}(y)\in O_{x}$. Either $x$ is not isolated, or $y=x=h^{n}(y)$ provided that $O_{x}$ is sufficiently tight. The latter case means $y$ is periodic, which is a contradiction with our assumption. ∎
We can now define the center of $X$.
Definition 5.8 (The center).
The maximal by inclusion compact invariant subset $X_{\mathrm{cent}}$ of $X$ satisfying the conclusion of Theorem 5.6 is called the center of $h$. As follows from the above, $X_{\mathrm{cent}}$ can be obtained as $X_{\alpha}$ for a sufficiently large ordinal $\alpha$, which also shows that $X_{\mathrm{cent}}$ with the listed properties exists.
Recall that, by the Cantor–Bendixon theorem, any compact metric space $X$ can be uniquely represented as a union of a perfect subset (possibly empty) and an at most countable subset. The former subset is called the perfect part $X_{\mathrm{perf}}$ of $X$, and the latter is called the countable part of $X$. By Theorem 5.6, if $X$ admits a self-homeomorphism without periodic points, then the perfect part of $X$ is nonempty; it contains the center $X_{\mathrm{cent}}$. We state this as a corollary.
Corollary 5.9.
If a compact metric space $X$ is at most countable, then any homeomorphism of $X$ with itself has a periodic point.
The operation of passing to the center behaves nicely under set-theoretic inclusion.
Lemma 5.10.
Suppose that $h:X\to X$ is a homeomorphism of a compact metric space $X$ without periodic points. For any compact $h$-invariant subset $Y\subset X$, one has $Y_{\mathrm{cent}}\subset X_{\mathrm{cent}}$.
Proof.
Consider the process leading to the formation of $X_{\mathrm{cent}}$ and $Y_{\mathrm{cent}}$. We will prove by transfinite induction that $Y_{\alpha}\subset X_{\alpha}$ for every ordinal $\alpha$. For $\alpha=0$, this is our assumption. Suppose now that $Y_{\alpha}\subset X_{\alpha}$ for some $\alpha$ and let us show the inclusion $Y_{\alpha+1}\subset X_{\alpha+1}$. By definition, $Y_{\alpha+1}$ consists of points $y\in Y_{\alpha}$, every neighborhood of which contains points of $Y_{\alpha}$ returning to this chosen neighborhood. Every such point is also a point of $X_{\alpha+1}$ by definition. It remains to note that the inclusion survives the operation of infinite intersection. ∎
5.3. Essential part
Now we combine the results of Sections 5.1 and 5.2. Let $X$ be a compact metric space, and $h:X\to X$ be a homeomorphism acting without periodic points. Also, consider a lower semicontinuous function $\xi:X\to\mathbb{R}_{>0}$ that is qusi-invariant under $h$. This will be a standing setup of Section 5.3.
Definition 5.11 (The essential part).
First, consider the $\xi$-accumulation set $X^{\xi}$ of $X$. Next, pass to the center $X_{*}$ of $X^{\xi}$, and call $X_{*}$ the essential part of $X$ with respect to $(h,\xi)$.
By Theorem 5.6, the essential part $X_{*}$ is a nonempty perfect set, whose points are nonwandering under the iterates of $h$. Consider a compact $h$-invariant subset of $Y$ and look at its essential part.
Theorem 5.12.
Let $h:X\to X$ be a homeomorphism of a compact metric space $X$ without periodic points, and $\xi:X\to\mathbb{R}_{>0}$ be a lower semicontinuous function quasi-invariant under $h$. If $Y\subset X$ is an $h$-invariant compact subspace, then $Y_{*}\subset X_{*}$, in particular $Y\cap X_{*}=\varnothing$ implies that $Y=\varnothing$.
Proof.
Let $X$ and $Y$ satisfy the assumptions of the theorem. By Lemma 5.5, one has $Y^{\xi}\subset X^{\xi}$. It now follows from Lemma 5.10 that $Y_{*}\subset X_{*}$. Recall that $Y_{*}$ is always nonempty, for a nonempty $Y$. Hence, if $Y\cap X_{*}=\varnothing$, then $Y=\varnothing$. ∎
6. Existence of aperiodic points
Say that a polygon exchange transformation $f$ is unbranched if $\Gamma_{1}\setminus\Gamma_{0}$ consists of several straight line intervals with pairwise disjoint closures, and the same holds for $\Gamma_{-1}\setminus\Gamma_{0}$. The main result of this section is the following theorem.
Theorem 6.1.
Let $f$ be an unbranched polygon exchange transformation such that, for at least one co-oriented line $L$, the dynamic Hadwiger invariant $\mathrm{Inv}_{L}(f)$ is nonzero. Then $f$ has aperiodic points.
Here, the assumption that $f$ is unbranched can probably be removed; however, it simplifies the arguments, and it does hold for polygon exchange transformations arising from outer billiards on regular polygons. Whether the assumption about a dynamic Hadwiger invariant being nonzero is essential, we do not know. Theorem 6.1, together with Corollary 3.7, imply Theorem 1.1, since the map $g_{X}$ from Section 3 is clearly unbranched and has non-vanishing dynamic Hadwiger invariants. Our standing assumption throughout the proof of Theorem 6.1 will be that $f$ is a non-periodic polygon exchange transformation, however, it has no aperiodic points.
6.1. Bounded levels
Recall that $\mathcal{J}$ is the ideal Julia set of $f$, and that $h_{f}:\mathcal{J}\to\mathcal{J}$ is the homeomorphism induced by $f$; it acts on $\mathcal{J}$ without periodic points. Since there are no aperiodic points of $f$, the bilevel function $\mathrm{bilev}$ is everywhere finite on $J$. Define the corresponding function $\xi:\mathcal{J}\to\mathbb{Z}_{\geqslant 0}$ as follows: the value of $\xi$ at an essential nest $(Q_{n})$ equals the maximum value of the function $\mathrm{bilev}$ on the impression of this nest. Observe that either the impression of $(Q_{n})$ is a single point, in which case there is only one value of $\mathrm{bilev}$, or the impression lies in $\Gamma_{k}\cup\Gamma_{-k}$ for some $k>0$, in which case the function $\mathrm{bilev}$ is bounded above by $k$. Here, we used our standing assumption that $f$ has no aperiodic points. Thus, in any case $\xi$ is well defined on $\mathcal{J}$. It is also easy to see that $\xi$ is lower-semicontinious: this follows from semi-continuity of $\mathrm{bilev}$, see Lemma 2.10.
Definition 6.2 (The essential part of $\mathcal{J}$).
Note that $\xi$ is quasi-invariant under $h_{f}$, more precisely, $\xi\circ h_{f}\leqslant\xi+1$ and $\xi\circ h^{-1}_{f}\leqslant\xi+1$. Hence, by Theorem 5.2, the subset $\mathcal{J}^{\xi}\subset\mathcal{J}$ is nonempty and $h_{f}$-invariant. Moreover, by definition of the $\xi$-accumulation set, the function $\xi$ is bounded on $\mathcal{J}^{\xi}$. Consider now the center $\mathcal{J}_{*}$ of $\mathcal{J}^{\xi}$. The subset $\mathcal{J}_{*}\subset\mathcal{J}^{\xi}$ is non-empty, perfect, $h_{f}$-invariant, and consists of nonwandering (under $h_{f}|_{\mathcal{J}_{*}}$) points. According to more general Definition 5.11, it is called the essential part of $\mathcal{J}$.
Denote by $J_{*}$ the union of the impressions of all nests from $\mathcal{J}_{*}$. Since the function $\xi$ is bounded on $\mathcal{J}_{*}$, there is some integer $n_{*}>0$ with the property that $J_{*}$ is contained in $\Gamma_{n_{*}}\cap\Gamma_{-n_{*}}$.
Lemma 6.3.
The set $J_{*}$ is perfect.
Proof.
First, let us show that $J_{*}$ is compact; it suffices to verify that $J_{*}$ is closed in $J$. Choose a sequence $z_{n}\in J_{*}$ converging to some $z\in J$. Also, for every $n$, let $(Q_{n,k})_{k}$ be an essential nest from $\mathcal{J}_{*}$ whose impression contains $z_{n}$. Define an essential nest $(Q_{k})$ so that each $Q_{k}$ coincides with $Q_{n,k}$ for infinitely many values of $n$. Passing to subsequences, we may assume that $Q_{k}=Q_{n,k}$ for all sufficiently large $n$. By definition of topology in $\mathcal{J}$, this means that the nests $(Q_{n,k})_{k}$ converge to $(Q_{k})$ as $n\to\infty$. Since $\mathcal{J}_{*}$ is compact, the limit nest $(Q_{k})$ lies in $\mathcal{J}_{*}$. Evidently, $z$ belongs to the impression of $(Q_{k})$, hence $z\in J_{*}$.
Consider any point $z\in J_{*}$ and let $(Q_{n})$ be a nest from $\mathcal{J}_{*}$ whose impression contains $z$. Since $\mathcal{J}_{*}$ is perfect and by Lemma 2.15, the compact polygon $\overline{Q}_{n}$ contains other points of $J_{*}$ for arbirtarily large $n$. Either the impression of $(Q_{n})$ is the singleton $\{z\}$, in which case the polygons $\overline{Q}_{n}$ lie in arbitrarily small neighborhoods of $z$, and it follows from the above that $z$ is not isolated in $J_{*}$, or the impression of $(Q_{n})$ is an infinite continuum, necessarily perfect, hence $z$ is not isolated in this case either. ∎
Two co-oriented lines $L$, $L^{\prime}$ are said to be parallel if $L$ and $L^{\prime}$ do not intersect or coincide, and the co-orientations of $L$ and $L^{\prime}$ match, that is, the positive half-plane with respect to $L$ is contained in that for $L^{\prime}$, or the other way around.
A non-periodic nest $(Q_{n})$ is said to be supported by a co-oriented line $L$ if, for sufficiently large $n$, some edge of $Q_{n}$ is contained in $L$, and $Q_{n}$ itself lies on the positive side of $L$. If a nest is not supported by a co-oriented line, then the impression of this nest consists of aperiodic points, as is clear from the definitions. Thus, our standing assumptions imply that every nest is supported by at least one co-oriented line. Given a co-oriented line $L$, say that a nest $(Q_{n})$ is an $L$-limit nest of $\mathcal{J}_{*}$ if there is a sequence of nests in $\mathcal{J}_{*}$ distinct from $(Q_{n})$ and converging to $(Q_{n})$ such that all nests of the sequence are supported by the same co-oriented line parallel to $L$.
Definition 6.4 (The $L$-parts of $\mathcal{J}_{*}$ and $J_{*}$).
Given a co-oriented line $L$, let $\mathcal{J}^{L}_{*}$ denote the set of all $L$-limit nests of $\mathcal{J}_{*}$; call it the $L$-part of $\mathcal{J}_{*}$. Finally, the $L$-part $J^{L}_{*}$ of $J_{*}$ is defined as the union of the impressions of all the nests from $\mathcal{J}^{L}_{*}$.
The set $J^{L}_{*}$ is closed, which is proved in the same way as the closedness of $J_{*}$ in Lemma 6.3. Since $\xi$ is bounded on $\mathcal{J}_{*}$, the set $J^{L}_{*}$ lies in a finite union of lines parallel to $L$. It follows also that $J^{L}_{*}$ is perfect: any point $z\in J^{L}_{*}$ is a limit of other points $z_{n}\in J_{*}$, and, except when $z$ is a vertex of $\Gamma_{n_{*}}\cup\Gamma_{-n_{*}}$, these $z_{n}$ cannot lie in $J^{L^{\prime}}_{*}$ for $L^{\prime}$ not parallel to $L$. Say that two lines $L_{1}$ and $L_{2}$ are transverse if they have a transverse intersection point, that is, if they are not parallel even if one disregards the co-orientations.
Lemma 6.5.
If co-oriented lines $L_{1}$, $L_{2}$ are transverse, then the corresponding sets $\mathcal{J}^{L_{1}}_{*}$ and $\mathcal{J}^{L_{2}}_{*}$ are disjoint. The set $\mathcal{J}_{*}$ is the union of $\mathcal{J}^{L}_{*}$ over all co-oriented lines $L$ containing edges of $\Gamma_{1}$. Similarly, $J_{*}$ is the union of $J^{L}_{*}$ over the same finite set of lines $L$.
We do not claim that $J^{L_{1}}_{*}$ and $J^{L_{2}}_{*}$ are disjoint when $L_{1}$ and $L_{2}$ are transverse; the claim of the lemma concerns only the sets $\mathcal{J}^{L_{1}}_{*}$ and $\mathcal{J}^{L_{2}}_{*}$.
Proof.
Suppose that $L_{1}$ and $L_{2}$ are transverse. First note that the intersection $J^{L_{1}}_{*}\cap J^{L_{2}}_{*}$ lies in the finite set formed by intersecting the finitely many lines containing $J^{L_{1}}_{*}$ with the finitely many lines containing $J^{L_{2}}_{*}$. On the other hand, the corresponding intersection $\mathcal{J}^{L_{1}}_{*}\cap\mathcal{J}^{L_{2}}_{*}$ is invariant under $h_{f}$, hence infinite if nonempty. Any point of $J^{L_{1}}_{*}\cap J^{L_{2}}_{*}$ can lie in the impressions of only finitely many nests, by Lemma 2.15. It follows that $J^{L_{1}}_{*}\cap J^{L_{2}}_{*}$ must also be infinite, a contradiction. Hence $\mathcal{J}^{L_{1}}_{*}$ and $\mathcal{J}^{L_{2}}_{*}$ are disjoint, as claimed.
Representing $\mathcal{J}_{*}$ as a union of $\mathcal{J}^{L}_{*}$ is straightforward: any nest of $\mathcal{J}_{*}$ is a limit point of $\mathcal{J}_{*}$, hence it is an $L$-limit for at least one $L$. The last claim of the lemma is immediate from the definitions. ∎
Choose a co-oriented line $L$ parallel to an edge of $\Gamma_{1}$. For each positive integer $n$, define the set $\Gamma_{n,L}$ as the union of all compact edges of $\Gamma_{n}$ that are parallel to $L$. We will refer to the direction of $L$ as the horizontal direction and to the positive side of $L$ as the upper side. Let $f_{L}$ denote the extension of $f$ from $\mathrm{dom}(f)$ to $\Gamma_{1,L}$ minus the vertices of $\Gamma_{1,L}$ so that $f_{L}$ is continuous from the upper side at all points of $\Gamma_{1,L}$ except the vertices. In other words, on the bottom horizontal edge of every level 1 piece (if any), the map $f_{L}$ acts by the same isometry as the restriction of $f$ to the given piece. Periodic, aperiodic, and boundary points of $f_{L}$ are defined in the same way as for $f$; the difference is that some periodic or aperiodic points of $f_{L}$ are boundary points of $f$.
Lemma 6.6.
The map $f_{L}$ is defined at all points of $J^{L}_{*}$ except possibly at finitely many vertices of $\Gamma_{1,L}$, and $J^{L}_{*}$ is $f_{L}$-invariant. Moreover, all points of $J^{L}_{*}$ are nonwandering under $f_{L}$.
Proof.
Any point $z\in J^{L}_{*}$ lies in the impression of some nest $(Q_{n})$ supported by a line through $z$ parallel to $L$. The image $(Q^{\prime}_{n})$ of this nest under $h_{f}$ contains the point $f_{L}(z)$ provided that $f_{L}$ is defined on $z$. It follows that $f_{L}(z)\in J^{L}_{*}$. Note, finally, that $f_{L}$ can be undefined at $z$ only if $z$ is one of the finitely many vertices of $\Gamma_{1,L}$. Finally, the fact that points of $J^{L}_{*}$ are nonwandering under the iterates of $f_{L}$ follows from the corresponding property of $h_{f}$. ∎
6.2. Geometry and dynamics of $J^{L}_{*}$
In this section, we discuss what the set $J^{L}_{*}$ may look like and how $f_{L}$ may act on this set. Assume that $J^{L}_{*}\neq\varnothing$ for a given choice of a co-oriented line $L$. We think of $L$ as being a horizontal line, and will refer to the positive side of $L$ as the upper side. Consider the union $I_{L}$ of all edges of $\Gamma_{1,L}$ that contain points of $J^{L}_{*}$ inside, i.e., not only as endpoints. Then $f_{L}$ acts on $I_{L}$ as a translation scissor congruence and takes $I_{L}$ into $\Gamma_{1,L}$. It can be extended to an interval exchange transformation $\varphi_{L}$ of $\Gamma_{1,L}$. Note that the dynamics of $\varphi_{L}$ and $f_{L}$ must match on $J^{L}_{*}$ but may be different otherwise. We need some general properties of interval exchange maps.
Theorem 6.7 ([KH95, NPT13]).
Let $\varphi$ be an interval exchange transformation of an interval (or a finite union of intervals) $I$. There exists a finite set of open $\varphi$-invariant subsets $I_{1}$, $\dots$, $I_{m}\subset I$ with the following properties:
  • Each $I_{i}$ is a finite union of open intervals;
  • Each $I_{i}$ is either a periodic component, i.e., $I_{i}$ is a cycle of intervals not properly contained in a cycle of bigger intervals, or a minimal component, i.e., all infinite forward or backward orbits in $I_{i}$ are dense;
The number $m$ is bounded above by the number of intervals being exchanged by $\varphi$.
See Theorem 14.5.13 in [KH95] and Theorem A of [NPT13] for a sharper upper bound on $m$. For our purposes, results of [KH95] are enough. By Lemma 6.6 and Theorem 6.7, the set $J^{L}_{*}$ consists of one or several minimal components of $\varphi_{L}$. In particular, it is a finite union of intervals on which $f_{L}$ acts as an interval exchange transformation.
Definition 6.8 (Cell space structures on $J^{L}_{*}$).
Endpoints of the components of $J^{L}_{*}\cap\mathrm{dom}(f_{L})$ are called forward 0-cells of $J^{L}_{*}$. Some of the 0-cells may lie inside components of $J^{L}_{*}$ rather than being their endpoints; such 0-cells are also vertices of $\Gamma_{1,L}$. Components of $J^{L}_{*}\cap\mathrm{dom}(f_{L})$ are called forward 1-cells of $J^{L}_{*}$. In this way, $J^{L}_{*}$ is equipped with a natural structure of a finite 1-dimensional cell space. Call this structure the forward cell structure. The backward cell structure is defined in the same way, replacing $f_{L}$ with $f_{L}^{-1}$. All forward and backward 1-cells of $J^{L}_{*}$ are straight line intervals parallel to $L$.
Recall that elementary sectors were introduced in Definition 2.16.
Definition 6.9 (Forward and backward trains).
Let $\varepsilon>0$ be a sufficiently small real number, in particular, smaller than the distance between any two distinct lines containing components of $J^{L}_{*}$. A forward trapezoid is defined as the trapezoid with horizontal bases such that the bottom base is a forward 1-cell of $J^{L}_{*}$, the height is $\varepsilon$, and, near the endpoints of its bottom base, the trapezoid coincides with elementary sectors. When $\varepsilon$ needs to be indicated, we speak of forward $\varepsilon$-trapezoids. Backward trapezoids are defined in the same way, replacing forward 1-cells with backward 1-cells. The union of the forward trapezoids is called the forward $\varepsilon$-train of $J_{L*}$, or just the forward train of $J_{L*}$ if $\varepsilon$ is fixed. Backward trains are defined similarly.
Denote the forward $\varepsilon$-train of $J^{L}_{*}$ by $\mathcal{T}^{L}_{*}(\varepsilon)$ or just by $\mathcal{T}^{L}_{*}$ if $\varepsilon$ is fixed. Evidently, $f(\mathcal{T}^{L}_{*})$ is then the corresponding backward train, Note that the union of the bottom bases of $\mathcal{T}^{L}_{*}$ is a subdivision of $J^{L}_{*}$, as well as the union of the bottom bases of $f(\mathcal{T}^{L}_{*})$. The following lemma is immediate from the definitions.
Lemma 6.10.
The map $f$ defines a translation scissors congruence from the forward train of $J^{L}_{*}$ to the backward train.
Note that the $\varepsilon$-trains of $J^{L}_{*}$, both forward and backward, are contained in an arbitrarily small neighborhood of $J^{L}_{*}$, provided that $\varepsilon$ is small enough for the chosen neighborhood. The set theoretic difference between the backward train and the forward train, with the same $\varepsilon$, consists of several triangles (possibly degenerate, i.e., line segments) cut out of the union of the two trains by $\Gamma_{1}$. Similarly, the forward train minus the backward train is a union of finitely many triangles cut out of the union of the two trains by $\Gamma_{-1}$.
Remark 6.11.
If $f$ is unbranched, then there can be no non-degenerate triangles cut out by $\Gamma_{1}$ or by $\Gamma_{-1}$. Therefore, the forward and the backward trains are subdivisions of the same polygon, and $f$ acts on this polygon as a polygon exchange transformation. Moreover, the restriction of this polygon exchange transformation to every component of its domain is a parallel translation by a vertor parallel to $L$, that is, by a horizontal vector.
In the proof of Theorem 6.1, the following property of the $\varepsilon$-trains will be needed.
Lemma 6.12.
Suppose that an essential nest $(Q_{n})$ belongs to $\mathcal{J}_{*}$, and fix $\varepsilon>0$. There is a co-oriented line $L$ parallel to an edge of $\Gamma_{1}$ and such that, for all sufficiently large $n$, the bi-piece $Q_{n}$ is contained in the $\varepsilon$-train of $J^{L}_{*}$.
Proof.
By Lemma 6.5, there is a unique (up to parallel translations) $L$ such that the nest $(Q_{n})$ belongs to $\mathcal{J}^{L}_{*}$. It follows, in particular, that $(Q_{n})$ is supported by $L$, possibly, after a parallel translation of $L$. Hence the impression of $(Q_{n})$ lies in $L$, and $Q_{n}$ is therefore contained in the $\varepsilon$-train of $J^{L}_{*}$ for all sufficiently large $n$. ∎
6.3. Horizontal polygon exchange transformations
Fix a co-oriented line $L$ and, as before, refer to its direction as the horizontal direction, and to its positive side as the upper side. Suppose that a polygon exchange transformation $f_{\Theta}$ of a polygon $\Theta$ translates all level 1 pieces by vectors parallel to $L$. In this case, $f_{\Theta}$ is called a horizontal polygon exchange transformation with respect to $L$. Our standing assumption in Section 6.3 is that $f_{\Theta}$ is a horizontal polygon exchange transformation with respect to $L$. There are several special properties of $f_{\Theta}$. Any horizontal line $L_{t}$ parallel to $L$ intersects $\Theta$ over a finite union of horizontal intervals. Note that $f_{\Theta}$ induces an interval exchange transformation of $\Theta\cap L_{t}$. Here, the parameter $t$ can be understood, e.g., as a height over $L=L_{0}$, and we write $\varphi_{t}$ as the interval exchange transformation of $\Theta\cap L_{t}$ induced by $f_{\Theta}$.
Theorem 6.13.
If $f_{\Theta}$ is a horizontal polygon exchange transformation with respect to $L$, then $\mathrm{Inv}_{L}(f_{\Theta})=0$.
Before giving a proof of Theorem 6.13, we make several preparatory remarks. There are finitely many lines $L_{t_{0}}$, $\dots$, $L_{t_{m}}$ such that the vertices of all level 1 pieces of $f_{\Theta}$ are contained in the union of these lines. Assume that the set of lines $L_{t_{i}}$ with the just mentioned property is chosen to be minimal by inclusion. In this case, the part $\Theta_{i}$ of $\Theta$ between $L_{t_{i}}$ and $L_{t_{i+1}}$ is a union of trapezoids with horizontal edges in $L_{t_{i}}$ and $L_{t_{i+1}}$. Some of these edges may degenerate to points, letting the corresponding trepezoids degenerate to triangles. Call the thus obtained trapezoids between $L_{t_{i}}$ and $L_{t_{i+1}}$ the trapezoids on the $i$-th floor. Observe that $f_{\Theta}$ acts as a horizontal polygon exchange transformation on $\Theta_{i}$. Knowing that the dynamic invariant $\mathrm{Inv}_{L}$ is additive with respect to polygonal subdivisions of the domain, it suffices to prove Theorem 6.13 for the union of all trapezoids on a given floor. Moreover, if we consider the first return map to any given trapezoid and then subdivide it further into floors if necessary, then we reduce Theorem 6.13 to the case when $\Theta$ is just an $L$-trapezoid, i.e., a trapezoid with top and bottom edges, possibly degenerate, parallel to $L$. We therefore assume now that $\Theta$ is an $L$-trapezoid.
Under the assumptions made, the dynamics of $f_{\Theta}$ reduces to that of interval exchange maps. Rescale the parameter $t$ so that the bottom and the top edges of $\Theta$ lie in $L_{0}$ and $L_{1}$, respectively. The family of interval exchange transformations $\varphi_{t}$, $t\in[0;1]$, is linear in the following sense: $\varphi_{t}$ translates the intervals $I_{1}(t)$, $\dots$, $I_{m}(t)$ by the same translation 1D vectors $\tau_{1}$, $\dots$, $\tau_{m}$ (independent of $t$), and the length $\lambda_{i}(t)$ of $I_{i}(t)$ depends linearly on $t\in[0;1]$ for all $i=1$, $\dots$, $m$. It is possible that $\lambda_{i}(0)=0$ or $\lambda_{i}(1)=0$ but $\lambda_{i}(t)>0$ for all $t\in(0;1)$. A family $\varphi_{t}$ of interval exchange transformations as just described will be referred to as a linear one parameter family of interval exchange transformations. Next, we recall some basic data associated with interval exchange transformations.
Definition 6.14 (Data associated with IETs).
Consider an interval exchange transformation $\varphi$ of some interval $I$. Write $I_{1}$, $\dots$, $I_{m}$ for the components of $\mathrm{dom}(\varphi)$. Assume that they are numbered from left to right. Let $\pi(i)$ be defined so that $\varphi(I_{i})$ is the $\pi(i)$-th component of $\mathrm{dom}(\varphi^{-1})$ from the left; then $\pi=\pi_{\varphi}$ is a permutation of $\{1,\dots,m\}$ called the permutation of $\varphi$. Define the vector of lengths of $\varphi$ as the vector $\lambda\in\mathbb{R}^{m}$ whose components are the lengths $\lambda_{1}$, $\dots$, $\lambda_{m}$ of subintervals $I_{1}$, $\dots$, $I_{m}$. The vector of translations of $\varphi$ is by definition the vector $\tau\in\mathbb{R}^{m}$ formed by the numbers $\tau_{1}$, $\dots$, $\tau_{m}$ such that $\varphi|_{I_{i}}$ acts as the translation $x\mapsto x+\tau_{i}$.
Recall also that the vector of lengths and a permutation $\pi$ of $\{1,\dots,m\}$ determine the vector of translations as follows: $\tau_{k}=\sum_{\pi(i)<\pi(k)}\lambda_{i}-\sum_{j<k}\lambda_{j}$. The latter linear relation can be rewritten in the matrix form as $\tau=\Omega\lambda$ and in coordinates as $\tau_{i}=\sum_{j=1}^{m}\Omega_{ij}\lambda_{j}$, where
$\Omega_{ij}=\begin{cases}1,&\hbox{if }\pi(i)>\pi(j)\hbox{ and }i<j,\\ -1,&\hbox{if }\pi(i)<\pi(j)\hbox{ and }i>j,\\ 0,&\hbox{in all other cases}.\end{cases}$
As is seen directly from the formula just displayed, the matrix $\Omega$ is skew symmetric. By the way, the skew-symmetry of $\Omega$ implies the (well-known) fact that SAF invariant of $\varphi$ takes values in $\bigwedge^{2}_{\mathbb{Q}}\mathbb{R}$, as the following brief computation demonstrates:
$\mathrm{Inv}(\varphi)=\sum_{i=1}^{m}\lambda_{i}\otimes\tau_{i}=\sum_{i,j=1}^{m}\Omega_{ij}\lambda_{i}\otimes\lambda_{j}=\sum_{i<j}\Omega_{ij}\lambda_{i}\wedge\lambda_{j},$
where $x\wedge y$ means $x\otimes y-y\otimes x$, and all tensor products are over $\mathbb{Q}$. Here, we used that $\lambda_{i}\otimes\Omega_{ij}\lambda_{j}=\Omega_{ij}\lambda_{i}\otimes\lambda_{j}$, since $\Omega_{ij}\in\mathbb{Q}$, and $\otimes$ is bilinear over $\mathbb{Q}$.
Lemma 6.15.
Let $(\varphi_{t})$ be a linear one parameter family of interval exchange transformations. Then the SAF invariant of $\varphi_{t}$ is independent of $t$.
Proof.
Write $\lambda_{i}(t)$ as $\lambda_{i}(0)+t\lambda^{\prime}_{i}(0)$ and substitute this expression into the formula $\sum\lambda_{i}(t)\otimes\tau_{i}$ for the invariant $\mathrm{Inv}(\varphi_{t})$. Recall that $\mathbb{R}\otimes\mathbb{R}$ can be viewed as a vector space over $\mathbb{R}$, with left multiplication by real scalars satisfying the identity $(ta)\otimes b=t(a\times b)$. It follows that $\mathrm{Inv}(\varphi_{t})$ can be written as the sum of two terms: the first term is independent of $t$ and is equal to $\mathrm{Inv}(\varphi_{0})$, whereas the second term is $t$ times the sum of $\lambda^{\prime}_{i}(0)\otimes\tau_{i}$ over $i$ ranging from 1 to $m$. Clearly, it suffices to show that this second term, equal to $\frac{d}{dt}\mathrm{Inv}(\varphi_{t})|_{t=0}$, vanishes (the derivative makes sense since $\varphi_{t}$ takes values in a two-dimensional $\mathbb{R}$-vector subspace of $\mathbb{R}\otimes\mathbb{R}$). Substituting the expression for $\tau$ as $\tau=\Omega\lambda(0)$, we obtain that
$\left.\frac{d}{dt}\mathrm{Inv}(\varphi_{t})\right|_{t=0}=\sum_{i,j=1}^{m}\Omega_{ij}\,\lambda^{\prime}_{i}(0)\otimes\lambda_{j}(0).$
Here, the factor $\Omega_{ij}$ has been moved to the front of the tensor product since $\Omega_{ij}\in\mathbb{Q}$.
On the other hand, differentiating each component of the matrix identity $\tau=\Omega\lambda(t)$ by $t$ and using that $\tau$ does not depend on $t$, we see that $\sum_{j}\Omega_{ij}\lambda^{\prime}_{j}(0)=0$. It remains to substitute the latter equality into the expression for $\frac{d}{dt}\mathrm{Inv}(\varphi_{t})|_{t=0}$ and use the skew symmetry of the matrix $\Omega$. We obtain that $\mathrm{Inv}(\varphi_{t})$ is constant with respect to $t$, as claimed. ∎
We are now ready to complete the proof of Theorem 6.13.
Proof of Theorem 6.13.
By the above, it suffices to assume that $\Theta$ is an $L$-trapezoid between $L_{0}$ and $L_{1}$, so that $\varphi_{t}$ for $t\in[0;1]$ form a linear one parameter family of interval exchange transformations. Since all the translation vectors are horizontal, the space $\mathbb{R}^{2}$ of translations reduces to $\mathbb{R}^{1}$, where all $\tau_{i}(t)$ lie. With this identification, the dynamic Hadwiger invariant of $f_{\Theta}$ identifies with the difference $\mathrm{Inv}(\varphi_{0})-\mathrm{Inv}(\varphi_{1})$ of two SAF invariants. By Lemma 6.15, the latter difference is zero. ∎
As it turns out, other dynamic Hadwiger invariants of $f_{\Theta}$ vanish as well.
Theorem 6.16.
All dynamic Hadwiger invariants of a horizontal polygon exchange transformation $f_{\Theta}$ are zero.
Proof.
Vanishing of $\mathrm{Inv}_{L}(f_{\Theta})$ follows from Theorem 6.13. Let $L^{\prime}$ be a line transverse to $L$; we want to show that $\mathrm{Inv}_{L^{\prime}}(f_{\Theta})=0$. It suffices to assume that $\Theta$ is an $L$-trapezoid between $L_{0}$ and $L_{1}$. Also, one may assume without loss of generality that the positive side of $L^{\prime}$ is on the right of $L^{\prime}$. Write $\ell^{\prime}$ for the length of the intersection of the strip between $L_{0}$ and $L_{1}$ with any line parallel to $L^{\prime}$. Suppose that $i_{1}<\dots<i_{k}$ are all the values of the index $i$, for which the left end of $I_{i}(t)$ for all $t\in(0;1)$ lies on an edge of $\Gamma_{1}(f_{\Theta})$ parallel to $L^{\prime}$. One also writes $j_{1}<\dots<j_{l}$ for the values of $j$ such that the right end of $I_{j}(t)$ for all $t\in(0;1)$ lies on an edge of $\Gamma_{1}(f_{\Theta})$ parallel to $L^{\prime}$. The value $\mathrm{Inv}_{L^{\prime}}(f_{\Theta})$ is by definition equal to
$\ell^{\prime}\otimes\left(\tau_{i_{1}}+\dots+\tau_{i_{k}}-\tau_{j_{1}}-\dots-\tau_{j_{l}}\right),$
and we need to show that the second factor equals zero.
Introduce an affine coordinate system in the plane so that one of the coordinate axes is horizontal, i.e., parallel to $L$. Denote by $x_{i}(t)$, respectively, $y_{i}(t)$, the horizontal coordinate of the left endpoint, respectively, the right endpoint, of $I_{i}(t)$. Similarly, write $x^{\prime}_{i}(t)$ and $y^{\prime}_{i}(t)$ for the corresponding horizontal coordinates of $\varphi_{t}(I_{i}(t))$. Note that
$x_{i_{1}}(t)+\dots+x_{i_{k}}(t)-y_{j_{1}}(t)-\dots-y_{j_{l}}(t)=$
=\ = $x^{\prime}_{i_{1}}(t)+\dots+x^{\prime}_{i_{k}}(t)-y^{\prime}_{j_{1}}(t)-\dots-y^{\prime}_{j_{l}}(t).$
Indeed, both in the left-hand side and in the right-hand side, terms corresponding to points of the edges of $\Gamma_{1}(f_{\Theta})$ parallel to $L^{\prime}$ inside $\Theta$ cancel out, and only those terms which correspond to the side edges of $\Theta$ survive. Finally, it remains to observe that $\tau_{i_{r}}=x^{\prime}_{i_{r}}(t)-x_{i_{r}}(t)$ and $\tau_{j_{s}}=y^{\prime}_{j_{s}}(t)-y_{j_{s}}(t)$, where $r=1$, $\dots$, $k$ and $s=1$, $\dots$, $l$. ∎
6.4. Proof of Theorem 6.1
Assume, as in Theorem 6.1, that $f$ is an unbranched polygon exchange transformation of a polygon $\Delta$ such that $\mathrm{Inv}_{L}(f)\neq 0$ for some co-oriented line $L$. By way of contradiction, we also assume that $f$ has no aperiodic points. Choose $\varepsilon>0$ so small that all $\varepsilon$-trains, associated with all possible co-oriented lines containing edges of $\Gamma_{1}$, are pairwise disjoint. Let $\mathcal{T}_{*}(\varepsilon)$ be the union of all these $\varepsilon$-trains, and write $\overline{\mathcal{T}}_{*}(\varepsilon)$ for its closure. Removing $\overline{\mathcal{T}}_{*}(\varepsilon)$ from $\Delta$ yields another (non-convex) polygon $\Delta_{\circ}$. Moreover, $\Delta_{\circ}$ is invariant under $f$; write $f_{\circ}$ for the restriction of $f$ to $\Delta_{\circ}$.
Lemma 6.17.
All the dynamic Hadwiger invariants of $f_{\circ}$ are the same as those of $f$. Hence, at least one of these invariants takes a nonzero value on $f_{\circ}$.
Proof.
Indeed, this follows from the additivity of the invariants and the fact that the restriction of $f$ to every train has all the dynamic Hadwiger invariants zero, by Theorem 6.16. ∎
Since $f_{\circ}$ has a nonzero dynamic Hadwiger invariant, it is not periodic. However, $f_{\circ}$ cannot have aperiodic points — all aperiodic points of $f_{\circ}$ would also be aperiodic points of $f$. The ideal Julia set $\mathcal{J}(f_{\circ})$ of $f_{\circ}$ and the essential part of it, $\mathcal{J}_{*}(f_{\circ})$, are therefore nonempty.
Lemma 6.18.
There is a natural topological embedding of $\mathcal{J}(f_{\circ})$ into $\mathcal{J}(f)$ that preserves the dynamics. In other words, $\mathcal{J}(f_{\circ})$ realizes as an $h_{f}$-invariant compact subset of $\mathcal{J}(f)$. Moreover, the latter subset is disjoint from $\mathcal{J}_{*}(f)$.
Proof.
Consider an essential nest $(Q^{\circ}_{n})$ of $f_{\circ}$. Evidently, each $Q^{\circ}_{n}$ is contained in a unique level $n$ bi-piece $Q_{n}$ of $f$. Thus, we obtain a nest $(Q_{n})$ of $f$. Moreover, the latter is essential since the former is. One obtains a set theoretic embedding of $\mathcal{J}(f_{\circ})$ into $\mathcal{J}(f)$, which is a homeomorphism to its image, by definition of the topologies in $\mathcal{J}(f_{\circ})$ and $\mathcal{J}(f)$. From now on, we identify $\mathcal{J}(f_{\circ})$ with its image in $\mathcal{J}(f)$. The $h_{f}$-invariance of $\mathcal{J}(f_{\circ})$ is clear. By Lemma 6.12, every nest $(Q_{n})$ that belongs to $\mathcal{J}_{*}(f)$ has the property that $Q_{n}$ is contained in a train of some $J^{L}_{*}$, for all sufficiently large $n$. It follows that $\mathcal{J}(f_{\circ})$ is disjoint from $\mathcal{J}_{*}(f)$. ∎
By Theorem 5.12, we obtain a contradiction with $\mathcal{J}(f_{\circ})$ being a nonempty compact $h_{f}$-invariant subset of $\mathcal{J}(f)$ disjoint from $\mathcal{J}_{*}(f)$, which proves Theorem 6.1.
7. Summary and perspectives
Our specific objective was to prove, for a regular $N$-gon $\Pi$ with $N\neq 3$, $4$, $6$, that the outer billiard map $f_{\Pi}$ has aperiodic points (Theorem 1.1). However, from a more general perspective, we also aimed at introducing tools for verifying aperiodicity of scissors automorphisms.
7.1. Proof of Theorem 1.1: a summary
Let $\Pi$ be a regular $N$-gon with $N\neq 3$, $4$, $5$, $6$. A theorem of Rukhovich (Theorem 3.1) allows to assume that $\Pi$ has an even number of vertices, equivalently, that $\Pi$ is centrally symmetric. By the results of Section 3.1, the polygonal domain $X$ defined there is invariant under $f_{\Pi}$; the restriction of $f_{\Pi}$ to $X$ is denoted by $f_{X}$. The central symmetry $M:z\mapsto-z$ of $\Pi$ is also a symmetry of $X$, therefore, $g_{X}:=M\circ f_{X}$ is a polygon exchange transformation of $X$. For $g_{X}$, it is not hard to compute the dynamic Hadwiger invariant (Theorem 3.6). Since this valuation does not vanish (Corollary 3.7), it follows (Theorem 2.29) that $g_{X}$ is not periodic. Further, Theorem 6.1 implies the existence of aperiodic points for $g_{X}$. Also, it is easy to see that $f_{X}$ has aperiodic points if and only if $g_{X}$ does — it suffices to note that $M$ commutes with $f_{X}$. In this way, we obtain the result of Theorem 1.1.
7.2. Comparison of the methods
In Sections 3 and 4, we gave two methods for establishing non-periodicity of certain scissors automorphisms. Method 1 uses dynamic Hadwiger invariants. For odd $N$, one can directly compute the dynamic Hadwiger invariant of $f_{X}^{2}$ (which acts by translations) — the only problem is that the computation gets technically more involved. An easier computation yields the value of the invariant for $g_{X}$, and, in fact, general properties of dynamic invariants imply that $\mathrm{Inv}_{L}(f_{X}^{2})=2\mathrm{Inv}_{L}(g_{X})$, for any choice of a co-oriented line $L$. Computing dynamic Hadwiger invariants for more general polygon exchange transformations may not be as easy. Note also that one needs to reduce a given piecewise isometry to a $\mathsf{T}$-scissors automorphisms in order to use Hadwiger invariants.
Method 2 was used to prove a special case of Theorem 3.8, namely, the case of a twice odd $N$. However, we believe that this method is also applicable to arbitrary regular polygons $\Pi$, possibly at the expense of more significant technical details. Similarly to method 1, the second method uses dynamic invariants, but this time in dimension one. The first step is therefore to reduce some properties of 2D dynamics to relevant properties of 1D dynamics — the notion of a boundary quasi-commutator and Theorem 4.3 serve this purpose. If a boundary quasi-commutator of a given map has nonzero SAF invariant, then the map cannot be periodic. Computations are simpler in the case when $N$ is twice odd and, for this reason, we worked out only this case (see Theorem 4.13); there are no principal obstructions to extending the computations to all $N$. From a general perspective, method 2 looks more robust: firstly, it applies to piecewise rotations, not only to piecewise translations, and secondly, it operates with SAF invariants of interval exchange transformations, which are better understood than their multi-dimensional analogs.
7.3. Possible extensions
In situation of Theorem 1.1, dynamic invariants can be computed explicitly. One possible direction for further research is extending Theorem 1.1 to all non-lattice quasi-rational polygons (a conjecture of Gutkin and Simanyi). An additional difficulty of this setup is that it is too general for relying on closed form formulas. General methods will be necessary for establishing the nonvanishing of dynamic invariants.
Related is a problem of better understanding what exactly nonzero dynamic invariants can say about dynamics, besides the lack of periodicity. This problem is not fully resolved even in dimension one, that is, for interval exchange transformations. It would be natural to first approach this problem either for special classes of interval exchange maps or for outer billiards on regular polygons.
7.4. Further questions
Some other problems are listed below:
  • Find interesting classes of multi-dimensional polytope exchange transformations, for which dynamic invariants are explicitly computable.
  • Relate polytope exchange transformations in dimension $d$ with flows on Euclidean $(d+1)$-manifolds; describe the role of dynamic invariants in terms of these flows.
  • Develop general methods for establishing (non)existence of dynamical self-similarity in the context of scissors automorphisms.